### v1

# Embodied Question Answering via Multi-LLM Systems

Abstract
Embodied Question Answering (EQA) is an important problem, which involves an agent exploring the environment to answer user queries. In the existing literature, EQA has exclusively been studied in single-agent scenarios, where exploration can be time-consuming and costly. In this work, we consider EQA in a multi-agent framework involving multiple large language models (LLM) based agents independently answering queries about a household environment. To generate one answer for each query, we use the individual responses to train a Central Answer Model (CAM) that aggregates responses for a robust answer. Using CAM, we observe a $50\%$ higher EQA accuracy when compared against aggregation methods for ensemble LLM, such as voting schemes and debates. CAM does not require any form of agent communication, alleviating it from the associated costs. We ablate CAM with various nonlinear (neural network, random forest, decision tree, XGBoost) and linear (logistic regression classifier, SVM) algorithms. Finally, we present a feature importance analysis for CAM via permutation feature importance (PFI), quantifying CAMs reliance on each independent agent and query context.
Embodied Question Answering via Multi-LLM Systems
Bhrij Patel University of Maryland, College Park                        Vishnu Sashank Dorbala University of Maryland, College Park Amrit Singh Bedi University of Central Florida                        Dinesh Manocha University of Maryland, College Park
1 Introduction
Embodied Question Answering (EQA) is an important task where an agent explores an environment, such as a household, to answer questions based on its observations. It has gained considerable interest in the past few years due to its applications for in-home robots and personalized assistants Das et al. (2018); Gordon et al. (2018); Duan et al. (2022a); Sima et al. (2023); Luo et al. (2023); Dorbala et al. (2024a); Zhu et al. (2023). For a seamless interface with humans, robotic agents must be equipped with natural language understanding Xu et al. (2021); Sun et al. (2024) and possess common-sense reasoning of the world. One form of common sense is having a relational understanding of household items and their usual place (i.e., a mug is highly likely to be in the kitchen or living room). Foundation models, specifically Large Language Models (LLMs) such as GPT Achiam et al. (2023) and LLama Touvron et al. (2023), have shown remarkable prowess in both natural language understanding and common sense reasoning. Recently, prior art has used these foundation models in EQA for query generation Dorbala et al. (2024a) and answering (Majumdar et al., 2024).
In existing works, only a single agent explores the environment with a limited field of view, inducing low coverage and high exploration costs. In contrast, for an LLM-based embodied agent to succeed in EQA, efficient exploration of the environment is needed as the questions will center around the state of the environment Das et al. (2018); Yu et al. (2019); Zhu et al. (2023); Majumdar et al. (2024). Prior work has studied the zero-shot capabilities of LLM-based agents to explore an unknown household environment Dorbala et al. (2023, 2024b); Majumdar et al. (2024).
The rise of Embodied Artificial Intelligence (EAI) Duan et al. (2022b) brings up the possibility of co-habited spaces with multiple agents and humans working in conjunction Puig et al. (2023). In this context, we consider a similar co-inhabited setup, and propose using a team comprising of multiple LLM-based agents or a Multi-LLM system for EQA. A Multi-LLM system allows for simultaneous exploration of the household environment in the same amount of time.
A key challenge in a multi-agent setup is with conflicting responses when agents are asked the same query. Since LLMs are blackboxes capable of generating free-form natural language, their responses to the same query may significantly vary Gal and Grosz (2022); Wang et al. (2024). In an embodied setting, such responses are further skewed by partial observations due to limited exploration. Prior work in ensemble LLM methods tackle this issue using weighted sum or voting schemes Li et al. (2024); Wei et al. (2022); Wang et al. (2022); Yang et al. (2024) and consensus-reaching debates between the LLM agents Chen et al. (2023a); Du et al. (2023); Xiong et al. (2023); Guo et al. (2024a); Chan et al. (2023). However, such methods can be vulnerable to scenarios where incorrect answers from malicious or poor-performing agents can influence the overall decision. An ideal EQA model would learn to identify incorrect agents and avoid being influenced by their responses. Furthermore, we would like to ideally avoid the communication costs incurred by debate-based approaches.
Main Results: To address the issues presented above, we solve EQA using a novel Multi-LLM agent approach. We instruct the agents to answer based on their observations a set of binary embodied questions about the household. We then train a Central Answer Model (CAM) on the answers of independent agents with supervised learning. The network directly outputs the final answer to a given question and does not require any communication between agents. Furthermore, our model learns which agents to rely on while producing its final answer, which reduces the end user’s effort in having to individually figure out the best agent response. Our novel contributions include:
  • Novel Multi-LLM EQA Framework: We propose a Central Answer Model (CAM) for Embodied Question Answering in a Multi-Agent setting. CAM is a classifier that aggregates multiple agent responses to predict an answer. We generate queries for various Matterport 3D Chang et al. (2017) environments. We train CAM on the labeled query dataset with multiple machine learning methods and show that our best model can achieve up to 50% higher accuracy than a majority vote and debate aggregation methods.
  • Integration with Exploration Systems: We evaluate our multi-agent framework on data gathered using an LLM-based exploration method Dorbala et al. (2023) on multiple agents in Matterport3D, demonstrating our system can work in conjunction with SOTA exploration methods in unknown environments.
  • Examining Agent Reliance: We present a feature importance analysis of CAM reliance on each independent agent using the metric, permutation feature importance (PFI) Breiman (2001); Fisher et al. (2018).
The rest of the paper is organized as follows: in Section 2, we provide a brief overview of the prior art relating to EQA and Multi-LLM systems; Section 3 gives a detailed overview of our approach; in Section 4, we describe our experimentation detail to evaluate our approach and follow with results showing the performance benefits compared to majority vote and consensus-reaching debate. Finally, in Section 7 and 8, we provide concluding remarks and a discussion on limitations, respectively.
Figure 1: Overview of our Multi-LLM framework. LLM-based agents that have explored the dynamic household are independently asked a binary “Yes/No” question and must answer based on their observations. Only one of the agents has the correct answer (“Yes”) while the other two agents are wrong due to either not exploring the relevant portions of the house enough or being malicious. The central network, however from supervised training, has learned to weigh the decision of the first agent more when given context about the item, room, and timestep.
Reference QA Exploration LLM-based Multi-Agent Feature Importance
(Majumdar et al., 2024)
(Singh et al., 2024)
Our Approach
Table 1: Table of our contributions and novelties compared to recent prior work. This work is the first to tackle Embodied EQA with LLM-based Multi-agent system and to provide a feature importance analysis on the framework.
2 Related Works
2.1 Embodied Question Answering
Embodied Question Answering (EQA) was first introduced by Das et al. (2018) and has become a popular task for robotics (Gordon et al., 2018, 2019; Duan et al., 2022a; Sima et al., 2023; Wijmans et al., 2019; Chen et al., 2019; Luo et al., 2019), where traditionally an agent is given a query such as “was there a plant in the living room?’ and explores its environment, in this case a house, to answer the question. Zhu et al. (2023) had alternated between an exploration phase and then a question phase and repeated the process until the query was answered. In our work, we first have an exploration gathering phase, then answer a set of questions. Further work has extended EQA to scenarios where the final answer revolves around multiple objects Yu et al. (2019), and where a target object is not explicitly stated in the question Tan et al. (2023). Tan et al. (2020) introduced multi-agent exploration and cooperation for question-answering in interactive environments. Dorbala et al. (2024a) provided a dataset of subjective or situational queries and was the first to use a generative approach for query creation, and Majumdar et al. (2024) designed the first open-vocabulary EQA dataset to test and benchmark foundation models on this task. Recently, Singh et al. (2024) studied the zero-shot LLM performance on visual QA with indoor house scenes. However, the LLM is given the scenes without any exploration of the environment. We evaluate our Multi-LLM approach with a SOTA LLM-based exploration method for observation gathering.
2.2 Multi-LLM Systems
Research in Multi-LLM systems have grown in popularity with the rise of foundation models for areas such as task planning Chen et al. (2023b), application development Wu et al. (2023), and chatbots Duan et al. (2023). Extensive studies have focused on their abilities to cooperate. Prior work has used multi-LLM debate to reach a consensus (Chan et al., 2023; Du et al., 2023; Guo et al., 2024a). Chen et al. (2023a) analyzes the consensus-seeking debate between multiple LLM agents with different personalities. For embodied AI, Guo et al. (2024b) have shown that multiple LLM-based agents can cooperate in teams by designating a leader for increased task efficiency. Other work have looked into using voting metrics of independent LLM-agents to arrive at a final answer (Li et al., 2024; Wei et al., 2022; Wang et al., 2022). Our work aims to remove the need for communication by training a central classifier on independent answers from multiple agents. Training a classifier in a supervised manner helps our approach become robust to unreliable agents that may influence overall answers in voting schemes.
Figure 2: Overview of experimentation pipeline. [TOP] We have an LLM answer a collection of embodied questions about the Matterport3D environment using the observations collected and stored by a specific agent during an exploration phase. The answers from the LLM are stored offline. [BOTTOM] In the central network training phase, we select multiple sets of answers from different agents to train the classifier with the ground-truth answer with supervised learning.
3 Our Approach: Multi-LLM System with Central Answer Network
In this work, we implement a two-step procedure that comprises a question-answering phase, and a central model training phase. The question-answering phases is agent-specific step while the classification-training step relies on combining the intermediate results of the question-answering phases from various agents. Before starting the question-answering phase, the agents have explored the environment to gather observations. Later, we see the performance of our approach when using a LLM-based SOTA method for exploration. We describe both phases in detail below.
3.1 Question-Answering Phase
In this phase, we generate a dataset of binary embodied questions specific to a Matterport 3D environment. Each LLM-based agent answers based on its independent observations from its exploration of the environment. To generate these questions we follow this template:
Do you think there is a [ITEM] in the [ROOM]? Use both common-sense reasoning about the object and room and the observation list given. Even if the relevant information is not in your observations, the answer can still be YES. Respond with YES or NO.
We then prompt the LLM to answer each question based on the stored observations of the agent. We pass the observations to the system prompt formatted as a dictionary, {ROOM NAME: [LIST OF ITEMS IN THE ROOM]}. The user prompt is a question following the template above and the system prompt is as follows:
You are an embodied agent that has explored a house. I prefer definite answers to questions which might require to explore. The observations are: [OBSERVATION DICTIONARY]
3.2 Central Model Training Phase
To give the final answer to each question, we train a model to give a final “Yes/No" output. The model input dimension is $2+|S|$, and the first two inputs account for the vectorized question in the form of “[object, room]”. The rest of the $|S|$ inputs come from entering the independent answers from each of the agents chosen for the multi-LLM system. During inference, we input unseen questions and agent answers into the network to output a test-time answer. Due to the modularity of our approach, for each set of independent responses, we train multiple machine learning models to produce multiple central answer models to compare against baseline aggregation approaches.
4 Experimentation
We now describe our experimentation in detail. With these experiments, we aim to 1) measure the accuracy of CAM for EQA and show its performance over other aggregation baselines (Section 4.2.2; 2) execute our Multi-LLM framework with observations gathered from a SOTA exploration method for unknown household environments, demonstrating its practicality for real-world scenarios (Section 5); and 3) provide a feature importance analysis to highlight the benefit of utilizing independent agent responses in our approach (Section 6) In all experiments we use GPT-4-turbo.
4.1 Query Generation
We extract the embodied question bank from the Matterport3D environments. For each scan, we extract an offline list of observed items detected in each image of the environment using the GLIP object detection algorithm Li et al. (2022).
For each item and room combination, we formulate the question using the template described in Section 3.1. Because these questions were extracted based on the observed items in the scan, we know the ground truth answer is “Yes".
To create questions with ground-truth answers “No", for each query generated with a “Yes" ground-truth, we formulate a question with the same item but with a random room that is not within the subset of rooms that contains that specific item.
4.2 EQA Performance Evaluation
In this section, to focus on the aggregation abilities of CAM, we have an oracle exploration where we explicitly say which rooms and items an agent has in its observations memory when answering the questions. In Section 5 we use a SOTA algorithm, LGX Dorbala et al. (2023) for the exploration phase.
Metric. To measure and compare the performances of our CAM approach against baseline aggregation approaches, we calculate the inference time accuracy, or correctness, of the model:
$\frac{1}{N}\sum_{i=1}^{N}\mathbbm{1}_{y_{i}=\hat{y}_{i}},$ (1)
where $y_{i}$ is the ground truth label of the test question, and $\hat{y}_{i}$ is the final answer prediction from CAM. We produce CAM models with various machine learning algorithms described in Table 2.
Algorithm Linearity
Neural Network (NN) Nonlinear
Random Forest (RF) Nonlinear
Decision Tree (DT) Nonlinear
XGBoost (XG) Nonlinear
RBF SVM (SVM) Nonlinear
Linear SVM (SVM-L) Linear
Logistic Regression (LR) Linear
Table 2: Table of different machine learning algorithms we implement to train CAM. We have a mix of linear and non-linear, parametric and non-parametric methods to thoroughly ablate the central model training phase.
4.2.1 Baselines
We compare these different variations of CAM against two other forms of the aggregation methods used for ensemble LLM methods. Neither method is learning-based and requires not training set. We describe them below:
  • Majority Vote (MV): For each test question, we take the majority vote of the independent answers from the agents of the Multi-LLM system. We do not directly take into account room or object with a majority vote
  • Debate: For each test question, we have the agents of the Multi-LLM system participate in a turn-based discussion where they relay their thoughts to one another for 2 rounds. To initial the debate we provide each LLM-based agent context with the following system prompt.
    System prompt: You are an embodied agent in a house. Your id is [INDEX]. Here are your observations from exploring the house: [OBSERVATION DICTIONARY]. Your initial answer to whether or not there was a [OBJECT] in the [ROOM] was [INITIAL ANSWER]. Other agents may have different answers. Please debate with the other agents to come to a consensus. Here is the conversation history: [CONVERSATION HISTORY]
    The last sentence is replaced with This is the beginning of the conversation. at the start of the debate when no conversation history is present. Below, we provide the user prompt to get the agent’s response for its turn during that round.
    User prompt: It is your turn to speak. Use your observation and conversation history to help.
    When prompted for the final answer at the end of the debate. The user prompt is Please give your final "Yes/No" answer. Give a definite answer.
We then collect the final answers from each agent and take the majority vote as the final prediction.
Figure 3: Outperforms Baselines; Test-time EQA accuracy measured by Equation 1 of CAM methods (blue) and baseline aggregation methods (orange). We report the mean and standard deviation over 5 seeds. The Multi-LLM system in these experiments uses independent answers from three agents that observations from bedroom, dining room, and kitchen. We use a 95%-5% train-test random split. We experiment on Matterport environments with [LEFT] 215 nodes and 15 distinct rooms [RIGHT] 53 nodes and 12 distinct rooms. Against the majority vote (MV) baseline, most CAM methods perform noticeably better during test time, with XGBoost (XG) achieving 50% higher accuracy. Against the debate baseline, the non-linear models perform at least as well, with XGBoost achieving 33% higher accuracy. The range in CAM methods performance highlights the need for careful model selection and tuning and that having ground truth labels does not guarantee significant improvement against the non-learning based baselines.
Figure 4: More Agents, Larger Test Set; With the $215$-node environment, we use a 80%-20% train-test random split, reporting the mean and standard deviation over 5 seeds. The Multi-LLM system in these experiments uses independent answers from $7$ agents. Each agent only used observations from one room to respond to queries. The $7$ rooms used that observations from the bedroom, dining room, living room, office, bathroom, family room and kitchen. Against the majority vote (MV) baseline, all CAM methods perform noticeably better during test time, with XGBoost (XG) achieving 46% higher accuracy.
Figure 5: LGX + CAM; Performance of CAM with $3$ independent agent that performed LGX for collecting observations. We use a 95%-5% train-test random split as in Figure 3. We experiment on Matterport environments with [LEFT] 215 nodes and [RIGHT] 53 nodes. Against both baselines, all CAM methods outperform during test-time inference, with XGBoost having the highest accuracy.
4.2.2 Results
We see the superior performance of CAM in Figure 3, where we run two sets of experiments, one to compare against MV and debating baseline in an $215$-node Matterport environment (left image) and the other in a $53$-node (right) environment. For each experiment, we report the mean and standard deviation accuracy calculated by Equation 1 over $5$ random seeds. We generate $600$ and $480$ queries for the $215$ and $53$-node environments, respectively, using the method described in Section 4.1. For both experiments, we train with $3$ independent answers to each query. The $3$ answers are from there LLM, where each time it answers it has only observations from one of the three rooms: bedroom, dining room, kitchen.
For the Central Model Training phase, the train-test random split of the generated queries and independent LLM-generated answers is $95$%-$5$%. In both environments, majority of the CAM method outperforms the MV and debating baselines. Specifically, XGBoost performs the best with $50$% $33$% higher accuracy than MV and debating, respectively. We also highlight the difference in time costs between the debate baseline and CAM. For each test question where three agents debate for $2$ rounds, the average inference time to make a final prediction is $30$ seconds, while the average time to train and run inference on the test set for XGBoost is $1.3$ seconds.
For a comparison on a larger test set with more input features using the $215$-node environment, in Figure 4 we have a random $80$%-$20$% train-test split. In this experiment, we have $7$ independent answers. Each answer is generated from the LLM, where the LLM only has access to the observations from a single room. The $7$ rooms are bedroom, kitchen, dining room, living room, family room, bathroom, and office. Each CAM method outperforms the baseline MV with XGBoost performing 46% higher.
We note that the two consistently lowest performing CAM methods and the ones closest to the MV and debating baselines are the linear models, Linear SVM and Logistic Regression Classifier. Intuitively, this similarity in performance makes sense as MV can be interpreted as a linear model with fixed equal weights, whereas Linear SVM and LR learn and adjust their weights based on the training data. Since we end up doing a majority vote on the final answers after the discussion, the debate baseline can also be construed as a linear model. These linear methods perform the worst suggesting the non-linearity of this EQA dataset.
5 CAM + Language-Guided Exploration
In this section, we evaluate our Multi-LLM framework by gathering observations using a SOTA exploration method. Rather than manually selecting which observations to provide to the LLM to answer the queries, we have all the LLM-based agents explore the household environment for a set amount of time before answering question.
In the exploration phase, each subordinate agent is set in the Matterport house scan graph and performs an augmented variation of Language-Guided Exploration (LGX) as proposed by Dorbala et al. (2023). In traditional LGX, the LLM agent observes the different immediate household items around itself and chooses which one to navigate towards to find a specific target item, forming a chain of “hops” to the item.
In our variation, we specifically provide the LLM with a system prompt that encourages the agent to explore as much of the house as possible. You are an embodied agent in a house. You want to aggressively explore the house and you want to find as many unique objects as you can.
Query Subject $\downarrow$ Object Kitchen Dining Room Bedroom Office
Kitchen 27.50 22.50 27.50 27.50 22.50
Dining Room 17.50 12.50 12.50 17.50 12.50
Bedroom 25.00 25.00 25.00 25.00 25.00
Office 12.50 5.00 5.00 12.50 7.50
Table 3: This table displays the feature importance of different decision tree-based CAM models. We see in cases where the queries only pertaining to the dining room (third row) that responses based only on dining room observation (fourth column) is not the most important feature. This insight suggests that CAM learned not to rely on that LLM agent.
The agent then travels from node to node in the house graph. For each timestep, we document the timestep, node, and items observed. When the agent reaches a node we can then access the stored results and add them to the list of observed items. We have the $3$ agents perform $10$ steps of this augmented LGX and store the observations. The rest of the experiment process follows the same setup as described in Section 3 with the Question-Answering and Central Model Training phases. We perform both these experiments in the same two environments as in Section 4.2.2 and using a $95\%$-$5\%$ random train-test split over 5 seeds.
The performance of the various CAM methods and baselines in the larger $215$-node graph is shown on the left image of Figure 5 and the performance for the $53$-node environment is shown on the right. From both plots, we again see CAM model outperforms both baselines as with Section 4.2.2. In fact, with LGX for observation collection, we see that on average, all CAM methods have a higher test time accuracy than the non-learning aggregation baselines.
Interestingly, in the 53-node environment debating baselines do slightly worse on average than MV. This degradation occurs due to agents who were initially correct being swayed by incorrect agents during conversation. Overall, our Multi-LLM system can be utilized in conjunction with an LLM-based exploration method, emphasizing its practicality in real-world use cases. In the next section, we provide a feature importance of CAM.
6 Feature Importance Analysis
In the context of EQA, feature importance is a metric by which input impacts the model’s final answer. For our Multi-LLM system, feature importance can be used to examine if any agent response or query context (i.e. room or object) is relied heavily upon for training the central answer model. To examine which features hold the most weight in the final decision of CAM model, we look into studying the variable, or feature importance Breiman (2001); Fisher et al. (2018), of the CAM model. We measure feature importance using permutation feature importance (PFI). We define the PFI the $i$-th feature as,
$PFI(i)=|a_{\text{val}}(m)-a_{\text{val}}^{i}(m)|,$ (2)
where $a_{\text{val}}(m)$ is the model $m$ accuracy measured by Equation 1 of the original validation set, and $a_{\text{val}}^{i}(m)$ is the accuracy with a modified validation set where the $i$-th feature column is randomly permuted. Intuitively, the more important a variable, the more the accuracy should change if the values are randomly shuffled because the variable is highly correlated with the target variable. In our Multi-LLM system for EQA, an agent that gives nearly-deterministic answers (e.g. “No” for almost all answers) would have a PFI close to 00.
To examine feature importance, we train multiple DT-based CAM models, each with a different filtered query dataset from the $215$-node environment. In each training, the features comprise of the room and object name to represent the query and $4$ independent responses from LLMs that had observations either from the kitchen, dining room, bedroom, or office. The dataset however will only consist of rows with a given “Room” value. The model is trained and tested, and we then calculate feature importance using Equation 2.
Table 3 summarizes our feature importance findings. The leftmost column specifies the room the dataset was filtered for, and the other columns show the average PFI of the object feature and independent responses for that filtered dataset over 5 trials. We see that in many cases, like when trained only on the dining room queries, the corresponding feature (i.e. response based only on dining room observation) does not have the highest PFI. We analyzed the offline LLM-generated answers based only on dining room observations and out of the $600$ responses, $549$ are “No”, or $91.5\%$. Thus, the CAM model has learned not to rely on these answers as much as the other responses even though they are based only on observations from the room in question. This outcome demonstrates how analyzing feature importance can help examine individual members of a multi-agent setup and further supports our approach utilizing independent agent responses.
7 Conclusion
In this paper, we tackle Embodied Question Answering with a novel Multi-LLM framework that aggregates the independent responses of LLM-based agents and produces a final answer. To the best of our knowledge, we are the first to use multiple LLM-based agents for EQA. We propose a central answer model (CAM) that learns to weigh the responses of each agent to decide on a single answer to the given query. We show that various machine learning methods are suitable to train CAM and consistently outperform non-learning-based aggregation methods for ensemble outputs, specifically majority voting and multi-LLM debating. A key benefit of our system is that it requires no communication during inference, reducing time costs. Furthermore, our approach can be combined with SOTA methods for exploration in unknown households. We end this work by providing a feature importance analysis to study the effect of individual agents and target object specification on the final answer. We hope that this work will encourage further research in Multi-LLM approaches for EQA as LLM-based embodied agents become more household.
8 Limitations and Future Work
This work has several limitations. First, the ground truth labels to train CAM can be difficult to obtain, especially if the household environment is constantly changing with non-stationary items. We leave it as future work to adapt this Multi-LLM setup for EQA in dynamic household environments. Second, our queries are all objective binary “Yes/No” questions. To better enhance practicality, future work should look into aggregation methods for situational, subjective, and non-binary questions. It would also be useful to remove the need to explicitly mention the target object in the query. Other areas directions for research could be then to apply our CAM aggregation method to question-answering tasks outside of Embodied AI such as long video understanding. Third, our results hinge on the quality of the observation data. In our work, we relied on the GLIP model. Future work should analyze the effect of scene understanding algorithms on our setup.

### v3

@@
 (2018).
-The rest of the paper is organized as follows: in Section 2, we provide a brief overview of the prior art relating to EQA and Multi-LLM systems; Section 3 gives a detailed overview of our approach; in Section 4, we describe our experimentation detail to evaluate our approach and follow with results showing the performance benefits compared to majority vote and consensus-reaching debate.
-Finally, in Section 7 and 8, we provide concluding remarks and a discussion on limitations, respectively.
+The rest of the paper is organized as follows: in Section 2, we provide a brief overview of the prior art relating to EQA and Multi-LLM systems; we formulate our EQA problem in Section 3.
+Section 4 gives a detailed overview of our approach; in Section 5, we describe our experimentation detail to evaluate our approach and follow with results showing the performance benefits compared to majority vote and consensus-reaching debate.
+Finally, in Section 8 and 9, we provide concluding remarks and a discussion on limitations, respectively.
 Figure 1: Overview of our Multi-LLM framework.
@@
 [BOTTOM] In the central network training phase, we select multiple sets of answers from different agents to train the classifier with the ground-truth answer with supervised learning.
+3 Problem Formulation
+In this section, we provide a mathematical formulation of our specific EQA setting to characterize the EQA problem we are tackling in this work.
+We consider a tuple of $(Q,S^{K})$, where $Q$ represents the set of queries, where each sample is a vector $\left[o,r,y\right]$ with object $o$ and room $r$ representing the question “Is there an $o$ in the $r$?”, and $y\in\{0,1\}$ is the binary ground-truth answer to the question.
+$S$ is the set of $K$ embodied agents that independently answer the questions.
+Let $s_{k}^{(o,r)}\in\{0,1\}$ be the independent answer to query $(o,r)$ of the $k$-th agent.
+Let $f$ be a function such that $f\left(o,r,\cup_{k=1}^{K}s_{i}^{(o,r)}\right)=\hat{y}\in\{0,1\}$.
+Given the query context $(o,r)$ and the independent agent answers to the context, $f$ should produce a final binary prediction to whether or not object $o$ is in room $r$.
 4 Our Approach: Multi-LLM System with Central Answer Network
@@
 Future work should analyze the effect of scene understanding algorithms on our setup.
+Appendix A Results on Additional Environments
+Below we provide additional results of our learning-based CAM methods against MV and debating on $7$ other Matterport environments.
+Similar to Figure 3, we use independent answers of three LLMs.
+One with observations only from dining room; another only with kitchen observations, and the last only with bedroom observations.
+Majority of the CAM methods consistently outperform the baseline majority vote (MV) and debate methods, with Decision Tree (DT) and XGBoost (XG) being the highest performing methods.
+Figure 6: 20 nodes, 5 distinct rooms
+Figure 7: 78 nodes, 11 distinct rooms
+Figure 8: 111 nodes, 14 distinct rooms
+Figure 9: 145 nodes, 15 distinct rooms
+Figure 10: 114 nodes, 16 distinct rooms
+Figure 11: 84 nodes, 13 distinct rooms
+Figure 12: 43 nodes, 5 distinct rooms
+Appendix B Decision Tree Visualization
+In this section, we showcase the potential interpretability of our learning-based CAM framework by visualizing the decision tree trained in Section 7.
+This tree in Figure 13 was trained on queries only pertaining to the kitchen and was trained on independent answers from four different LLM agents.
+One agent had only observations from kitchen; another only dining room; another bedroom; and the last one only had office observations.
+We see that the main feature used was actually the object query.
+We provide the object encoding labels in Table 4.
+Figure 13: Visualization of Decision Tree from Section 7.
+Object Label Encoding $o$
+appliance 0
+armchair 1
+bathtub 2
+bed 3
+board 4
+bookcase 5
+books 6
+cabinet 7
+chair 8
+clothes 9
+counter 10
+curtain 11
+cushion 12
+desk 13
+dresser 14
+dumbells 15
+hairbrush 16
+headphones 17
+jumprope 18
+mirror 19
+mug 20
+nightstand 21
+phone 22
+picture 23
+plant 24
+playing cards 25
+refrigerator 26
+sink 27
+sofa 28
+table 29
+television 30
+toilet 31
+toothbrush 32
+towel 33
+tv remote 34
+wallet 35
+water bottle 36
+whiteboard 37
+window 38
+wristwatch 39
+Table 4: Mapping of object labels to encodings $o$

### v4

@@
-Embodied Question Answering via Multi-LLM Systems
+Central Answer Modeling for an Embodied Multi-LLM System
 Abstract
@@
 To generate one answer for each query, we use the individual responses to train a Central Answer Model (CAM) that aggregates responses for a robust answer.
-Using CAM, we observe a $50\%$ higher EQA accuracy when compared against aggregation methods for ensemble LLM, such as voting schemes and debates.
+While prior Question Answering (QA) work has used a central module based on answers from multiple LLM-based experts, we specifically look at applying this framework to embodied LLM-based agents that must physically explore the environment first to become experts on their given environment to answer questions.
+Our work is the first to utilize a central answer model framework with embodied agents that must rely on exploring an unknown environment.
+We set up a variation of EQA where instead of the agents exploring the environment after the question is asked, the agents first explore the environment for a set amount of time and then answer a set of queries.
+Using CAM, we observe a $46\%$ higher EQA accuracy when compared against aggregation methods for ensemble LLM, such as voting schemes and debates.
 CAM does not require any form of agent communication, alleviating it from the associated costs.
 We ablate CAM with various nonlinear (neural network, random forest, decision tree, XGBoost) and linear (logistic regression classifier, SVM) algorithms.
-Finally, we present a feature importance analysis for CAM via permutation feature importance (PFI), quantifying CAMs reliance on each independent agent and query context.
-Embodied Question Answering via Multi-LLM Systems
-Bhrij Patel University of Maryland, College Park                        Vishnu Sashank Dorbala University of Maryland, College Park Amrit Singh Bedi University of Central Florida                        Dinesh Manocha University of Maryland, College Park
-1 Introduction
+We experiment in various topological graph environments and examine the case where one of the agents is malicious and purposes contributes responses it believes to be wrong.
+Finally, we present a feature importance analysis for CAM via permutation feature importance (PFI).
+We provide and code and query dataset for further research.
+Introduction
 Embodied Question Answering (EQA) is an important task where an agent explores an environment, such as a household, to answer questions based on its observations.
-It has gained considerable interest in the past few years due to its applications for in-home robots and personalized assistants Das et al.
-(2018); Gordon et al.
-(2018); Duan et al.
-(2022a); Sima et al.
-(2023); Luo et al.
-(2023); Dorbala et al.
-(2024a); Zhu et al.
-(2023).
-For a seamless interface with humans, robotic agents must be equipped with natural language understanding Xu et al.
-(2021); Sun et al.
-(2024) and possess common-sense reasoning of the world.
+It has gained considerable interest in the past few years due to its applications for in-home robots and personalized assistants (Das et al.
+2018; Gordon et al.
+2018; Duan et al.
+2022a; Sima et al.
+2023; Luo et al.
+2023; Dorbala et al.
+2024a; Zhu et al.
+2023).
+For a seamless interface with humans, robotic agents must be equipped with natural language understanding (Xu et al.
+2021; Sun et al.
+2024) and possess common-sense reasoning of the world.
 One form of common sense is having a relational understanding of household items and their usual place (i.e., a mug is highly likely to be in the kitchen or living room).
-Foundation models, specifically Large Language Models (LLMs) such as GPT Achiam et al.
-(2023) and LLama Touvron et al.
-(2023), have shown remarkable prowess in both natural language understanding and common sense reasoning.
-Recently, prior art has used these foundation models in EQA for query generation Dorbala et al.
-(2024a) and answering (Majumdar et al., 2024).
+Foundation models, specifically Large Language Models (LLMs) such as GPT (Achiam et al.
+2023) and LLama (Touvron et al.
+2023), have shown remarkable prowess in both natural language understanding and common sense reasoning.
+Recently, prior art has used these foundation models in EQA for query generation (Dorbala et al.
+2024a) and answering (Majumdar et al.
+2024).
 In existing works, only a single agent explores the environment with a limited field of view, inducing low coverage and high exploration costs.
-In contrast, for an LLM-based embodied agent to succeed in EQA, efficient exploration of the environment is needed as the questions will center around the state of the environment Das et al.
-(2018); Yu et al.
-(2019); Zhu et al.
-(2023); Majumdar et al.
-(2024).
-Prior work has studied the zero-shot capabilities of LLM-based agents to explore an unknown household environment Dorbala et al.
-(2023, 2024b); Majumdar et al.
-(2024).
-The rise of Embodied Artificial Intelligence (EAI) Duan et al.
-(2022b) brings up the possibility of co-habited spaces with multiple agents and humans working in conjunction Puig et al.
-(2023).
+In contrast, for an LLM-based embodied agent to succeed in EQA, efficient exploration of the environment is needed as the questions will center around the state of the environment (Das et al.
+2018; Yu et al.
+2019; Zhu et al.
+2023; Majumdar et al.
+2024).
+Prior work has studied the zero-shot capabilities of LLM-based agents to explore an unknown household environment (Dorbala, Mullen Jr, and Manocha 2023; Dorbala et al.
+2024b; Majumdar et al.
+2024).
+The rise of Embodied Artificial Intelligence (EAI) (Duan et al.
+2022b) brings up the possibility of co-habited spaces with multiple agents and humans working in conjunction (Puig et al.
+2023).
 In this context, we consider a similar co-inhabited setup, and propose using a team comprising of multiple LLM-based agents or a Multi-LLM system for EQA.
-A Multi-LLM system allows for simultaneous exploration of the household environment in the same amount of time.
+A Multi-LLM system allows for simultaneous exploration of the environment in the same amount of time.
 A key challenge in a multi-agent setup is with conflicting responses when agents are asked the same query.
-Since LLMs are blackboxes capable of generating free-form natural language, their responses to the same query may significantly vary Gal and Grosz (2022); Wang et al.
-(2024).
+Since LLMs are blackboxes capable of generating free-form natural language, their responses to the same query may significantly vary (Gal and Grosz 2022; Wang et al.
+2024).
 In an embodied setting, such responses are further skewed by partial observations due to limited exploration.
-Prior work in ensemble LLM methods tackle this issue using weighted sum or voting schemes Li et al.
-(2024); Wei et al.
-(2022); Wang et al.
-(2022); Yang et al.
-(2024) and consensus-reaching debates between the LLM agents Chen et al.
-(2023a); Du et al.
-(2023); Xiong et al.
-(2023); Guo et al.
-(2024a); Chan et al.
-(2023).
-However, such methods can be vulnerable to scenarios where incorrect answers from malicious or poor-performing agents can influence the overall decision.
+Prior work in ensemble LLM methods tackle this issue using weighted sum or voting schemes (Li et al.
+2024; Wei et al.
+2022; Wang et al.
+2022; Yang et al.
+2024) and consensus-reaching debates between the LLM agents (Chen et al.
+2023a; Du et al.
+2023; Xiong et al.
+2023; Guo et al.
+2024a; Chan et al.
+2023; Li, Patel, and Du 2023).
+However, such methods can be vulnerable to scenarios where incorrect answers from malicious or poor-performing agents can influence the overall decision (Amayuelas et al.
+2024).
+Figure 1: Overview of our Multi-LLM framework.
+LLM-based agents that have explored the dynamic household are independently asked a binary “Yes/No” question and must answer based on their observations.
+Only one of the agents has the correct answer (“Yes”) while the other two agents are wrong due to either not exploring the relevant portions of the house enough or being malicious.
+The central network, however from supervised training, has learned to weigh the decision of the first agent more when given context about the item, room, and timestep.
 An ideal EQA model would learn to identify incorrect agents and avoid being influenced by their responses.
 Furthermore, we would like to ideally avoid the communication costs incurred by debate-based approaches.
-Main Results: To address the issues presented above, we solve EQA using a novel Multi-LLM agent approach.
+Main Results: To address the issues presented above, we apply a Multi-LLM system to solve this embodied task.
 We instruct the agents to answer based on their observations a set of binary embodied questions about the household.
@@
 Our novel contributions include:
-  • Novel Multi-LLM EQA Framework: We propose a Central Answer Model (CAM) for Embodied Question Answering in a Multi-Agent setting.
+  • Multi-LLM QA Framework with Embodied Navigation: We train a Central Answer Model (CAM) for Question Answering in an Embodied Multi-Agent setting.
+We depart from prior EQA setups where the agents explore after being asked a question.
+Rather we have the agents explore before being asked a set of question on their environment.
 CAM is a classifier that aggregates multiple agent responses to predict an answer.
-We generate queries for various Matterport 3D Chang et al.
-(2017) environments.
-We train CAM on the labeled query dataset with multiple machine learning methods and show that our best model can achieve up to 50% higher accuracy than a majority vote and debate aggregation methods.
-  • Integration with Exploration Systems: We evaluate our multi-agent framework on data gathered using an LLM-based exploration method Dorbala et al.
-(2023) on multiple agents in Matterport3D, demonstrating our system can work in conjunction with SOTA exploration methods in unknown environments.
-  • Examining Agent Reliance: We present a feature importance analysis of CAM reliance on each independent agent using the metric, permutation feature importance (PFI) Breiman (2001); Fisher et al.
-(2018).
-The rest of the paper is organized as follows: in Section 2, we provide a brief overview of the prior art relating to EQA and Multi-LLM systems; we formulate our EQA problem in Section 3.
-Section 4 gives a detailed overview of our approach; in Section 5, we describe our experimentation detail to evaluate our approach and follow with results showing the performance benefits compared to majority vote and consensus-reaching debate.
-Finally, in Section 8 and 9, we provide concluding remarks and a discussion on limitations, respectively.
-Figure 1: Overview of our Multi-LLM framework.
-LLM-based agents that have explored the dynamic household are independently asked a binary “Yes/No” question and must answer based on their observations.
-Only one of the agents has the correct answer (“Yes”) while the other two agents are wrong due to either not exploring the relevant portions of the house enough or being malicious.
-The central network, however from supervised training, has learned to weigh the decision of the first agent more when given context about the item, room, and timestep.
-Reference QA Exploration LLM-based Multi-Agent Feature Importance
-(Majumdar et al., 2024)
-(Singh et al., 2024)
+We generate queries for various Matterport 3D (Chang et al.
+2017) environments.
+We train CAM on the labeled query dataset with multiple machine learning methods and show that our best model can achieve up to 46% higher accuracy than a majority vote and debate aggregation methods.
+While prior work has used a central answer module to output a final answer from a mixture of answers from multiple LLM experts (Puerto, Şahin, and Gurevych 2023; Si et al.
+2023; Jiang, Ren, and Lin 2023), we are the first to apply the framework to embodied agents that must first explore an environment.
+  • Integration with Exploration Systems: We evaluate our multi-agent framework on data gathered using an LLM-based exploration method (Dorbala, Mullen Jr, and Manocha 2023) on multiple agents in Matterport3D, demonstrating our system can work in conjunction with SOTA exploration methods in unknown environments.
+  • Examining Agent and Feature Reliance: We present a feature importance analysis of CAM reliance on each independent agent using the metric, permutation feature importance (PFI) (Breiman 2001; Fisher, Rudin, and Dominici 2018).
+We also look at the agreement between the individual agent and the final answer of the central answer model and the baselines.
+Similarly to Si et al.
+(2023), some of our methods for the central answer model are interpretable and provide a Decision Tree visualization of one our models.
+Reference Exploration Central Module Multi-Agent Feature Importance
+(Majumdar et al. 2024)
+(Puerto, Şahin, and Gurevych 2023)
 Our Approach
@@
 This work is the first to tackle Embodied EQA with LLM-based Multi-agent system and to provide a feature importance analysis on the framework.
-2 Related Works
-2.1 Embodied Question Answering
+Related Works
+Embodied Question Answering and Exploration
 Embodied Question Answering (EQA) was first introduced by Das et al.
-(2018) and has become a popular task for robotics (Gordon et al., 2018, 2019; Duan et al., 2022a; Sima et al., 2023; Wijmans et al., 2019; Chen et al., 2019; Luo et al., 2019), where traditionally an agent is given a query such as “was there a plant in the living room?’ and explores its environment, in this case a house, to answer the question.
+(2018) and has become a popular task for robotics (Gordon et al.
+2018; Gordon, Fox, and Farhadi 2019; Duan et al.
+2022a; Sima et al.
+2023; Wijmans et al.
+2019; Chen et al.
+2019; Luo et al.
+2019), where traditionally an agent is given a query such as “was there a plant in the living room?’ and explores its environment, in this case a house, to answer the question.
 Zhu et al.
 (2023) had alternated between an exploration phase and then a question phase and repeated the process until the query was answered.
-In our work, we first have an exploration gathering phase, then answer a set of questions.
-Further work has extended EQA to scenarios where the final answer revolves around multiple objects Yu et al.
-(2019), and where a target object is not explicitly stated in the question Tan et al.
-(2023).
+Further work has extended EQA to scenarios where the final answer revolves around multiple objects (Yu et al.
+2019), and where a target object is not explicitly stated in the question (Tan et al.
+2023).
 Tan et al.
@@
 (2024) designed the first open-vocabulary EQA dataset to test and benchmark foundation models on this task.
-Recently, Singh et al.
-(2024) studied the zero-shot LLM performance on visual QA with indoor house scenes.
+Recently, Singh, Pavlakos, and Stamoulis (2024) studied the zero-shot LLM performance on visual QA with indoor house scenes.
 However, the LLM is given the scenes without any exploration of the environment.
-We evaluate our Multi-LLM approach with a SOTA LLM-based exploration method for observation gathering.
-2.2 Multi-LLM Systems
-Research in Multi-LLM systems have grown in popularity with the rise of foundation models for areas such as task planning Chen et al.
-(2023b), application development Wu et al.
-(2023), and chatbots Duan et al.
-(2023).
+We evaluate our Multi-LLM approach with a SOTA LLM-based exploration method for observation gathering introduced by Dorbala, Mullen Jr, and Manocha (2023).
+Note: In our work, we deviate from the traditional EQA approach.
+We first have an exploration gathering phase, then answer a set of questions on the same environment.
+We do so for two reasons: 1) to provide a setup to better compare against Multi-LLM QA baselines: majority voting and debating; 2) we believe this setup is better suited for the scenario where the user asks a question about the items in their homes and quickly needs an answer.
+Multi-LLM Systems and Experts
+Research in Multi-LLM systems have grown in popularity with the rise of foundation models for areas such as task planning (Chen et al.
+2023b), application development (Wu et al.
+2023), and chatbots (Duan et al.
+2023).
 Extensive studies have focused on their abilities to cooperate.
-Prior work has used multi-LLM debate to reach a consensus (Chan et al., 2023; Du et al., 2023; Guo et al., 2024a).
+Prior work has used multi-LLM debate to reach a consensus (Chan et al.
+2023; Du et al.
+2023; Guo et al.
+2024a; Chen, Saha, and Bansal 2023).
 Chen et al.
-(2023a) analyzes the consensus-seeking debate between multiple LLM agents with different personalities.
+(2023a) analyzes the consensus-seeking debate between multiple LLM agents with different personalities, and Amayuelas et al.
+(2024) investigates the susceptibility of a multi-LLM discussion to an adversarial agent that purposefully tries to lead an incorrect answer choice.
 For embodied AI, Guo et al.
 (2024b) have shown that multiple LLM-based agents can cooperate in teams by designating a leader for increased task efficiency.
-Other work have looked into using voting metrics of independent LLM-agents to arrive at a final answer (Li et al., 2024; Wei et al., 2022; Wang et al., 2022).
+Other work have looked into using voting metrics of independent LLM-agents to arrive at a final answer (Li et al.
+2024; Wei et al.
+2022; Wang et al.
+2022).
 Our work aims to remove the need for communication by training a central classifier on independent answers from multiple agents.
 Training a classifier in a supervised manner helps our approach become robust to unreliable agents that may influence overall answers in voting schemes.
+A central architecture to output a final answer from multiple answers from different LLM experts has been explored previously (Puerto, Şahin, and Gurevych 2023; Si et al.
+2023; Jiang, Ren, and Lin 2023).
+Our approach differs from these prior works as the various LLM agents we utilize become experts in the environments from an embodied exploration phase and we show it can be used with off-the-shelf exploration methods.
+We believe that our approach can be considered a variation of Mixture-of-Experts (MoE) (Eigen, Ranzato, and Sutskever 2013), as early MoE methods relied on training ML models like SVM to learn the weightings of different experts, as our CAM framework does.
+Our main difference is that the MoE module also takes into account answers from embodied agents explore the environment.
 Figure 2: Overview of experimentation pipeline.
-[TOP] We have an LLM answer a collection of embodied questions about the Matterport3D environment using the observations collected and stored by a specific agent during an exploration phase.
+[LEFT] We have an LLM answer a collection of embodied questions about the Matterport3D environment using the observations collected and stored by a specific agent during an exploration phase.
 The answers from the LLM are stored offline.
-[BOTTOM] In the central network training phase, we select multiple sets of answers from different agents to train the classifier with the ground-truth answer with supervised learning.
-3 Problem Formulation
+[RIGHT] In the central network training phase, we select multiple sets of answers from different agents to train the classifier with the ground-truth answer with supervised learning.
+Multi-LLM System with Central Answer Model (CAM) for EQA
+Problem Formulation
 In this section, we provide a mathematical formulation of our specific EQA setting to characterize the EQA problem we are tackling in this work.
@@
 Let $s_{k}^{(o,r)}\in\{0,1\}$ be the independent answer to query $(o,r)$ of the $k$-th agent.
-Let $f$ be a function such that $f\left(o,r,\cup_{k=1}^{K}s_{i}^{(o,r)}\right)=\hat{y}\in\{0,1\}$.
+Let $f$ be a function such that $f\left(o,r,\cup_{k=1}^{K}s_{k}^{(o,r)}\right)=\hat{y}\in\{0,1\}$.
 Given the query context $(o,r)$ and the independent agent answers to the context, $f$ should produce a final binary prediction to whether or not object $o$ is in room $r$.
-4 Our Approach: Multi-LLM System with Central Answer Network
-In this work, we implement a two-step procedure that comprises a question-answering phase, and a central model training phase.
-The question-answering phases is agent-specific step while the classification-training step relies on combining the intermediate results of the question-answering phases from various agents.
-Before starting the question-answering phase, the agents have explored the environment to gather observations.
-Later, we see the performance of our approach when using a LLM-based SOTA method for exploration.
-We describe both phases in detail below.
-4.1 Question-Answering Phase
+In this work, we present CAM as a viable method of the function $f$ over baselines majority voting and debating.
+We now describe CAM in the next section.
+Our Approach
+We implement a two-step procedure that comprises a question-answering phase, and a central model training phase.
+The question-answering phase is an agent-specific step while the classification-training step relies on combining the intermediate results of the question-answering phases from various agents.
+Before starting the question-answering phase, the agents explored the environment to gather observations.
+Later, we examine the performance of our approach when using a LLM-based SOTA method for exploration.
+We describe both phases below.
+Question-Answering Phase
 In this phase, we generate a dataset of binary embodied questions specific to a Matterport 3D environment.
-Each LLM-based agent answers based on its independent observations from its exploration of the environment.
+Matterport3D is a large-scale dataset popular for scene-understanding tasks and embodied AI research, containing RGB-D panoramic views of various building environments like homes.
+The data is stored as a graph for each environment where the nodes are different physical locations within the home environment that an embodied agent can occupy.
+We use a topological graph environment because we are only concerned with high-level exploration policy in the house environment.
+Our approach can be used with other similar environments such as HM3D (Ramakrishnan et al.
+2021).
+Each LLM-based agent’s answers are based on independent observations from its exploration of the environment.
+To emphasize, the agents differ in their collected observations, leading to differing answers of the same query.
 To generate these questions we follow this template:
@@
 We pass the observations to the system prompt formatted as a dictionary, {ROOM NAME: [LIST OF ITEMS IN THE ROOM]}.
-The user prompt is a question following the template above and the system prompt is as follows:
-You are an embodied agent that has explored a house. I prefer definite answers to questions which might require to explore. The observations are: [OBSERVATION DICTIONARY]
-4.2 Central Model Training Phase
-To give the final answer to each question, we train a model to give a final “Yes/No" output.
-The model input dimension is $2+|S|$, and the first two inputs account for the vectorized question in the form of “[object, room]”.
-The rest of the $|S|$ inputs come from entering the independent answers from each of the agents chosen for the multi-LLM system.
+The user prompt is a question following the template above and the system prompt is as follows: You are an embodied agent that has explored a house.
+I prefer definite answers to questions.
+The observations are: [OBSERVATION DICTIONARY]
+Central Model Training Phase
+To give the final answer to each question, we train a model $f$ to give a final, binary “Yes/No” output.
+The model input dimension is $2+K$, where $K$ is the number of agents.
+The first two inputs account for the vectorized question in the form of “[object, room]”, $(o,r)$.
+The rest of the $K$ inputs come from entering the independent answers from each of the agents chosen for the multi-LLM system, $\cup_{k=1}^{K}s_{k}^{(o,r)}$.
+$f$ then outputs a predicted answer $\hat{y}$ and is then optimized with the ground truth answer $y$.
 During inference, we input unseen questions and agent answers into the network to output a test-time answer.
 Due to the modularity of our approach, for each set of independent responses, we train multiple machine learning models to produce multiple central answer models to compare against baseline aggregation approaches.
-5 Experimentation
+Note: The classifier is overfitted to a given environment because the Multi-LLM embodied system is only for a specific environment.
+With our EQA variation, the agent must answer questions on what it has explored in the given environment.
+This overfitting allows for the personalization of our systems to the household environment.
+In our potential use case of a user asking a question of whether or not an object is located in a certain area of their environment, the agents should be trained and asked in the same environment.
+Experimentation
 We now describe our experimentation in detail.
-With these experiments, we aim to 1) measure the accuracy of CAM for EQA and show its performance over other aggregation baselines (Section 5.2.2; 2) execute our Multi-LLM framework with observations gathered from a SOTA exploration method for unknown household environments, demonstrating its practicality for real-world scenarios (Section 6); and 3) provide a feature importance analysis to highlight the benefit of utilizing independent agent responses in our approach (Section 7) In all experiments we use GPT-4-turbo.
-5.1 Query Generation
+With these experiments, we aim to 1) measure the accuracy of CAM for EQA and show its performance over other aggregation baselines (Results section); 2) and 3) provide a feature importance analysis to highlight the benefit of utilizing independent agent responses in our approach (Feature Importance Analysis Section) In all experiments we use GPT-4-turbo.
+Query Generation
 We extract the embodied question bank from the Matterport3D environments.
-For each scan, we extract an offline list of observed items detected in each image of the environment using the GLIP object detection algorithm Li et al.
-(2022).
-For each item and room combination, we formulate the question using the template described in Section 4.1.
-Because these questions were extracted based on the observed items in the scan, we know the ground truth answer is “Yes".
-To create questions with ground-truth answers “No", for each query generated with a “Yes" ground-truth, we formulate a question with the same item but with a random room that is not within the subset of rooms that contains that specific item.
-5.2 EQA Performance Evaluation
-In this section, to focus on the aggregation abilities of CAM, we have an oracle exploration where we explicitly say which rooms and items an agent has in its observations memory when answering the questions.
-In Section 6 we use a SOTA algorithm, LGX Dorbala et al.
-(2023) for the exploration phase.
+For each scan, we extract an offline list of observed items detected in each image of the environment using the GLIP object detection algorithm (Li et al.
+2022).
+The items space observed by GLIP can be found in the Appendix in Table 6.
+For each item and room combination, we formulate the question using the template described in Section Question-Answering Phase.
+Because these questions were extracted based on the observed items in the scan, we know the ground truth answer is “Yes”.
+To create questions with ground-truth answers “No”, for each query generated with a “Yes” ground-truth, we formulate a question with the same item but with a random room that is not within the subset of rooms that contains that specific item.
+Note: Our setup is highly dependent on the object detector used.
+Error can thus be introduced by the object detection with GLIP and then the LLM answer generation.
+Some outputs of GLIP included toilets in the living room and a TV in the hallway.
+When the LLM agent has a (room, item) in its observation memory in the system prompt, it will usually say Yes if asked about said room and item, but when it does not have it in memory, the questions are harder to answer, leading to error in answers generated by the independent LLM agent.
+More error analysis on agent answers will be studied in further in the Feature Importance Analysis section.
+Evaluation
+We use a SOTA algorithm, LGX (Dorbala, Mullen Jr, and Manocha 2023) for the exploration phase.
 Metric. To measure and compare the performances of our CAM approach against baseline aggregation approaches, we calculate the inference time accuracy, or correctness, of the model:
-$\frac{1}{N}\sum_{i=1}^{N}\mathbbm{1}_{y_{i}=\hat{y}_{i}},$ (1)
+$\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}_{y_{i}=\hat{y}_{i}},$ (1)
 where $y_{i}$ is the ground truth label of the test question, and $\hat{y}_{i}$ is the final answer prediction from CAM.
-We produce CAM models with various machine learning algorithms described in Table 2.
-Algorithm Linearity
-Neural Network (NN) Nonlinear
-Random Forest (RF) Nonlinear
-Decision Tree (DT) Nonlinear
-XGBoost (XG) Nonlinear
-RBF SVM (SVM) Nonlinear
-Linear SVM (SVM-L) Linear
-Logistic Regression (LR) Linear
-Table 2: Table of different machine learning algorithms we implement to train CAM.
-We have a mix of linear and non-linear, parametric and non-parametric methods to thoroughly ablate the central model training phase.
-5.2.1 Baselines
+We produce CAM models with various machine learning algorithms: Neural Network (NN), Random Forest (RF), Decision Tree (DT), XGBoost (XG), RBF SVM (SVM), Linear SVM (SVM-L), Logistic Regression (LR).
+The NN was implemented with Pytorch using 3 linear layers.
+The dimension were 16, 8, 1 in that order with a Sigmoid activation function at the end to ensure the output was in between 0 and 1.
+XGBoost used the default settings from the xgboost library package.
+RF from the sklearn library used $1000$ trees, and all other methods used the default settings from sklearn.
+Baselines
 We compare these different variations of CAM against two other forms of the aggregation methods used for ensemble LLM methods.
-Neither method is learning-based and requires not training set.
+Neither method is learning-based and requires no training set.
 We describe them below:
@@
     User prompt: It is your turn to speak. Use your observation and conversation history to help.
-    When prompted for the final answer at the end of the debate. The user prompt is Please give your final "Yes/No" answer. Give a definite answer.
+    When prompted for the final answer at the end of the debate. The user prompt is Please give your final ”Yes/No” answer. Give a definite answer.
 We then collect the final answers from each agent and take the majority vote as the final prediction.
-Figure 3: Outperforms Baselines; Test-time EQA accuracy measured by Equation 1 of CAM methods (blue) and baseline aggregation methods (orange).
-We report the mean and standard deviation over 5 seeds.
-The Multi-LLM system in these experiments uses independent answers from three agents that observations from bedroom, dining room, and kitchen.
-We use a 95%-5% train-test random split.
-We experiment on Matterport environments with [LEFT] 215 nodes and 15 distinct rooms [RIGHT] 53 nodes and 12 distinct rooms.
-Against the majority vote (MV) baseline, most CAM methods perform noticeably better during test time, with XGBoost (XG) achieving 50% higher accuracy.
-Against the debate baseline, the non-linear models perform at least as well, with XGBoost achieving 33% higher accuracy.
-The range in CAM methods performance highlights the need for careful model selection and tuning and that having ground truth labels does not guarantee significant improvement against the non-learning based baselines.
-Figure 4: More Agents, Larger Test Set; With the $215$-node environment, we use a 80%-20% train-test random split, reporting the mean and standard deviation over 5 seeds.
-The Multi-LLM system in these experiments uses independent answers from $7$ agents.
-Each agent only used observations from one room to respond to queries.
-The $7$ rooms used that observations from the bedroom, dining room, living room, office, bathroom, family room and kitchen.
-Against the majority vote (MV) baseline, all CAM methods perform noticeably better during test time, with XGBoost (XG) achieving 46% higher accuracy.
+Language-Guided Exploration
 Figure 3: LGX + CAM; Performance of CAM with $3$ independent agent that performed LGX for collecting observations.
-We use a 95%-5% train-test random split as in Figure 3.
-We experiment on Matterport environments with [LEFT] 215 nodes and [RIGHT] 53 nodes.
+We use a 90%-10% train-test random split.
+We experiment on a Matterport environment with 215 nodes.
+We report the mean and standard deviation over $5$ trials.
 Against both baselines, all CAM methods outperform during test-time inference, with XGBoost having the highest accuracy.
-5.2.2 Results
-We see the superior performance of CAM in Figure 3, where we run two sets of experiments, one to compare against MV and debating baseline in an $215$-node Matterport environment (left image) and the other in a $53$-node (right) environment.
-For each experiment, we report the mean and standard deviation accuracy calculated by Equation 1 over $5$ random seeds.
-We generate $600$ and $480$ queries for the $215$ and $53$-node environments, respectively, using the method described in Section 5.1.
-For both experiments, we train with $3$ independent answers to each query.
-The $3$ answers are from there LLM, where each time it answers it has only observations from one of the three rooms: bedroom, dining room, kitchen.
-For the Central Model Training phase, the train-test random split of the generated queries and independent LLM-generated answers is $95$%-$5$%.
-In both environments, majority of the CAM method outperforms the MV and debating baselines.
-Specifically, XGBoost performs the best with $50$% $33$% higher accuracy than MV and debating, respectively.
-We also highlight the difference in time costs between the debate baseline and CAM.
-For each test question where three agents debate for $2$ rounds, the average inference time to make a final prediction is $30$ seconds, while the average time to train and run inference on the test set for XGBoost is $1.3$ seconds.
-For a comparison on a larger test set with more input features using the $215$-node environment, in Figure 4 we have a random $80$%-$20$% train-test split.
-In this experiment, we have $7$ independent answers.
-Each answer is generated from the LLM, where the LLM only has access to the observations from a single room.
-The $7$ rooms are bedroom, kitchen, dining room, living room, family room, bathroom, and office.
-Each CAM method outperforms the baseline MV with XGBoost performing 46% higher.
-We note that the two consistently lowest performing CAM methods and the ones closest to the MV and debating baselines are the linear models, Linear SVM and Logistic Regression Classifier.
-Intuitively, this similarity in performance makes sense as MV can be interpreted as a linear model with fixed equal weights, whereas Linear SVM and LR learn and adjust their weights based on the training data.
-Since we end up doing a majority vote on the final answers after the discussion, the debate baseline can also be construed as a linear model.
-These linear methods perform the worst suggesting the non-linearity of this EQA dataset.
-6 CAM + Language-Guided Exploration
-In this section, we evaluate our Multi-LLM framework by gathering observations using a SOTA exploration method.
-Rather than manually selecting which observations to provide to the LLM to answer the queries, we have all the LLM-based agents explore the household environment for a set amount of time before answering question.
-In the exploration phase, each subordinate agent is set in the Matterport house scan graph and performs an augmented variation of Language-Guided Exploration (LGX) as proposed by Dorbala et al.
-(2023).
-In traditional LGX, the LLM agent observes the different immediate household items around itself and chooses which one to navigate towards to find a specific target item, forming a chain of “hops” to the item.
-In our variation, we specifically provide the LLM with a system prompt that encourages the agent to explore as much of the house as possible.
+Accuracy (%)
+CAM-based Methods (Ours) Baselines
+Environment Size (num of nodes) NN RF DT XG SVM SVM-L LR MV Debate
+20 65.56 63.33 73.33 73.33 64.44 43.33 53.33 58.89 58.89
+43 61.00 63.00 86.00 82.00 62.00 60.00 64.00 41.00 40.00
+53 50.83 68.33 78.75 81.25 64.17 50.83 50.83 44.58 44.17
+58 50.00 74.00 84.00 87.00 74.00 55.00 53.00 54.00 53.00
+78 65.00 70.45 85.45 86.36 70.45 57.27 55.91 42.27 41.82
+114 64.06 77.19 81.25 87.50 70.94 55.63 60.31 51.88 47.19
+145 65.33 75.33 83.00 87.67 69.00 58.33 59.33 46.67 45.33
+215 66.33 72.67 79.67 88.33 69.67 56.00 56.33 42.00 43.00
+Accuracy (%) with Malicious Agent
+CAM-based Methods (Ours) Baselines
+Environment Size (num of nodes) NN RF DT XG SVM SVM-L LR MV Debate
+20 61.11 62.22 73.33 73.33 64.44 44.44 53.33 61.11 58.89
+53 60.42 68.33 78.75 81.25 64.17 49.58 50.83 46.67 44.17
+145 68.00 75.33 83.00 87.67 69.00 59.33 59.33 47.67 45.67
+215 64.33 73.67 79.67 88.33 69.67 56.33 56.33 43.67 43.00
+Table 2: Performance measured in accuracy (Eq 1) of CAM with $3$ independent agent that performed LGX for collecting observations.
+We use a 90%-10% train-test random split.
+We experiment in $8$ different environments over $5$ trials.
+Against both baselines, all CAM methods outperform during test-time inference, with XGBoost having the highest accuracy.
+We repeated the experiment in $4$ of the environments but turned one of the agent malicious by providing responses opposite to what it believes are the correct answers.
+In cases where one agent is malicious, our CAM method is still better suited than MV and debate.
+We have all the LLM-based agents explore the household environment for a set amount of time before answering questions.
+In the exploration phase, each subordinate agent is set in the Matterport house scan graph and performs an augmented variation of Language-Guided Exploration (LGX) as proposed by Dorbala, Mullen Jr, and Manocha (2023).
+In traditional LGX, the LLM agent observes the immediate household items around itself and chooses which one to navigate to find a specific target item, forming a chain of “hops” to the item.
+In our variation, we instead provide the LLM with a system prompt that encourages the agent to explore as much of the house as possible.
 You are an embodied agent in a house.
 You want to aggressively explore the house and you want to find as many unique objects as you can.
+The agent then travels from node to node in the house graph.
+For each timestep, we document the timestep, node, and items observed.
+When the agent reaches a node we can then access the stored results and add them to the list of observed items.
+We have the $3$ agents perform $10$ steps of this augmented LGX and store the observations.
+The rest of the experiment process follows the same setup as described in the Approach Section with the Question-Answering and Central Model Training phases.
+We use a $90\%$-$10\%$ random train-test split over $5$ seeds (0 to 4 are used as the seeds).
+Experiments are ran on an Apple M1 Pro and macOS 14.5.
+Results: The performance of the various CAM methods and baselines in the $8$ graph environments is shown in Table 2.
+We test in environments of different sizes (measured in number of nodes) to showcase the generalization of the CAM approach across different size complexities.
+From the table, we see CAM model consistently outperforms both baselines in accuracy (Eq 1.
+We report the average accuracy over $5$ seeds.
+We see that the non-linear methods consistently achieve higher accuracy than the linear methods, suggesting that the dataset contains no hyperplane separating the queries with “Yes” and “No” ground-truth answers.
+Specifically, XGBoost attains the highest accuracy in all but one environment, suggesting that the boosted ensemble method can best fit the generated query-answer data.
+Interestingly, in all but one of the environments, debating does slightly worse on average than MV.
+This degradation occurs due to agents who were initially correct being swayed by incorrect agents during conversation.
+Overall, our Multi-LLM system can be utilized in conjunction with an LLM-based exploration method, emphasizing its practicality in real-world use cases.
+Method Agent 1 Agent 2 Agent 3
+XG 38 $\pm$ 8.3 38 $\pm$ 7.0 38 $\pm$ 6.2
+MV 97 $\pm$ 0.1 95 $\pm$ 0.1 96 $\pm$ 1.6
+debate 97 $\pm$ 1.4 93 $\pm$ 3.1 91 $\pm$ 3.6
+Table 3: Agreement (%) between each LGX agent and the final answers of 1) XGBoost-based CAM (XG), 2) Majority Vote (MV) and 3) debate in the $215$-node environment.
+We report the average and standard deviation over the $5$ trials.
+We see that XGBoost learns not to agree with any of the agents to achieve its high accuracy, suggesting more exploration or better object detection is needed.
+Agreement: We analyze the agreement between each agent and the XGBoost model (the best performing CAM-based method), majority vote, and debate.
+We measure agreement between an agent and each central module method by the percentage of similar answers:
+$Agr(s_{k},f)=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}_{s_{k}^{i}=\hat{y}_{i}},$ (2)
+where $i$ is index of the query parameterized by $(o,r)$.
+Table 3 reports the agreement over $5$ trials and we see that XGBoost has learned not to agree with the agents as much as MV and debate to attain its high accuracy.
+This low agreement could stem from the errors introduced by the object detection of GLIP and the question-answering of each individual agent.
+We further analyze the independent answers impact on training in the Feature Importance Analysis Section.
 Query Subject $\downarrow$ Object Kitchen Dining Room Bedroom Office
@@
 This insight suggests that CAM learned not to rely on that LLM agent.
-The agent then travels from node to node in the house graph.
-For each timestep, we document the timestep, node, and items observed.
-When the agent reaches a node we can then access the stored results and add them to the list of observed items.
-We have the $3$ agents perform $10$ steps of this augmented LGX and store the observations.
-The rest of the experiment process follows the same setup as described in Section 4 with the Question-Answering and Central Model Training phases.
-We perform both these experiments in the same two environments as in Section 5.2.2 and using a $95\%$-$5\%$ random train-test split over 5 seeds.
-The performance of the various CAM methods and baselines in the larger $215$-node graph is shown on the left image of Figure 5 and the performance for the $53$-node environment is shown on the right.
-From both plots, we again see CAM model outperforms both baselines as with Section 5.2.2.
-In fact, with LGX for observation collection, we see that on average, all CAM methods have a higher test time accuracy than the non-learning aggregation baselines.
-Interestingly, in the 53-node environment debating baselines do slightly worse on average than MV.
-This degradation occurs due to agents who were initially correct being swayed by incorrect agents during conversation.
-Overall, our Multi-LLM system can be utilized in conjunction with an LLM-based exploration method, emphasizing its practicality in real-world use cases.
+Malicious Agent: We also show the benefits of CAM where one of the three agents is malicious.
+To turn the agent malicious we invert its independent answers so that all “Yes” answers turn to “No” and vice versa.
+We then train all CAM methods and execute baselines regularly.
+Thus, the malicious agent contributes an answer it believes to be incorrect.
+From Table 2, we see that our CAM methods still outperform the MV and debate baselines.
+We see that the debate baseline again does worse the majority vote, showing that the incorrect agent was able to sway the overall decision that was initially going to be correct.
 In the next section, we provide a feature importance of CAM.
-7 Feature Importance Analysis
-In the context of EQA, feature importance is a metric by which input impacts the model’s final answer.
+Feature Importance Analysis
+In the context of machine learning, feature importance is a metric by which input impacts the model’s final answer.
 For our Multi-LLM system, feature importance can be used to examine if any agent response or query context (i.e.
 room or object) is relied heavily upon for training the central answer model.
-To examine which features hold the most weight in the final decision of CAM model, we look into studying the variable, or feature importance Breiman (2001); Fisher et al.
-(2018), of the CAM model.
+To examine which features hold the most weight in the final decision of CAM model, we look into studying the variable, or feature importance (Breiman 2001; Fisher, Rudin, and Dominici 2018), of the CAM model.
 We measure feature importance using permutation feature importance (PFI).
@@
 To examine feature importance, we train multiple DT-based CAM models, each with a different filtered query dataset from the $215$-node environment.
-In each training, the features comprise of the room and object name to represent the query and $4$ independent responses from LLMs that had observations either from the kitchen, dining room, bedroom, or office.
+To better highlight and focus on the feature importance aspect, we have an oracle exploration where we explicitly say which rooms and items an agent has in its observations memory when answering the questions.
+This feature importance analysis could be applied to dynamic agents that move around with LGX.
+In each training, the features comprise the room and object name to represent the query and $4$ independent responses from LLMs that had observations either from the kitchen, dining room, bedroom, or office.
 The dataset however will only consist of rows with a given “Room” value.
@@
 Thus, the CAM model has learned not to rely on these answers as much as the other responses even though they are based only on observations from the room in question.
-This outcome demonstrates how analyzing feature importance can help examine individual members of a multi-agent setup and further supports our approach utilizing independent agent responses.
-8 Conclusion
-In this paper, we tackle Embodied Question Answering with a novel Multi-LLM framework that aggregates the independent responses of LLM-based agents and produces a final answer.
-To the best of our knowledge, we are the first to use multiple LLM-based agents for EQA.
+This outcome demonstrates how analyzing feature importance can help examine individual members of a multi-agent setup for error analysis and further supports our approach utilizing independent responses.
+Please see the Appendix for a visualization of the Decision Tree to highlight the benefits of the interpretable ML-based CAM methods.
+Conclusion and Limitations
+In this paper, we tackle a variation of Embodied Question Answering with a Multi-LLM framework that aggregates the independent responses of LLM-based agents and produces a final answer.
+While other methods have used a central module to aggregate independent agent answers for QA tasks, we specifically investigate this method in the context of embodied exploration in a household environment.
 We propose a central answer model (CAM) that learns to weigh the responses of each agent to decide on a single answer to the given query.
@@
 Furthermore, our approach can be combined with SOTA methods for exploration in unknown households.
+We also show that CAM is more robust to malicious agents than the majority vote and debate baslines.
 We end this work by providing a feature importance analysis to study the effect of individual agents and target object specification on the final answer.
-We hope that this work will encourage further research in Multi-LLM approaches for EQA as LLM-based embodied agents become more household.
-9 Limitations and Future Work
+We hope that this work will encourage research in embodied Multi-LLM for QA as LLM-based agents become more household.
 This work has several limitations.
@@
 To better enhance practicality, future work should look into aggregation methods for situational, subjective, and non-binary questions.
-It would also be useful to remove the need to explicitly mention the target object in the query.
-Other areas directions for research could be then to apply our CAM aggregation method to question-answering tasks outside of Embodied AI such as long video understanding.
-Third, our results hinge on the quality of the observation data.
-In our work, we relied on the GLIP model.
-Future work should analyze the effect of scene understanding algorithms on our setup.
-Appendix A Results on Additional Environments
-Below we provide additional results of our learning-based CAM methods against MV and debating on $7$ other Matterport environments.
-Similar to Figure 3, we use independent answers of three LLMs.
-One with observations only from dining room; another only with kitchen observations, and the last only with bedroom observations.
-Majority of the CAM methods consistently outperform the baseline majority vote (MV) and debate methods, with Decision Tree (DT) and XGBoost (XG) being the highest performing methods.
-Figure 6: 20 nodes, 5 distinct rooms
-Figure 7: 78 nodes, 11 distinct rooms
-Figure 8: 111 nodes, 14 distinct rooms
-Figure 9: 145 nodes, 15 distinct rooms
-Figure 10: 114 nodes, 16 distinct rooms
-Figure 11: 84 nodes, 13 distinct rooms
-Figure 12: 43 nodes, 5 distinct rooms
-Appendix B Decision Tree Visualization
-In this section, we showcase the potential interpretability of our learning-based CAM framework by visualizing the decision tree trained in Section 7.
+Appendix A Appendix
+Appendix B Code and Dataset
+We provide an anonymous link to our code  [^1] https://anonymous.4open.science/r/CAM-embodied-llm-54CE/. The query dataset extracted is also within the submitted supplementary materials folder.
+Appendix C Multi-LLM vs LLM
+To motivate multiple LLM agents we provide a comparison between 1 agent and 3 agents in Table 5.
+Environment Size Accuracy (%) of 1 Agent Accuracy (%) of 3 Agents (MV)
+$53$ nodes $41.67$ $42.00$
+$215$ node $44.17$ $44.58$
+Table 5: Rise of accuracy with multiple dynamic agents
+Appendix D Additional Visualization of Results
+In this section, to focus on the aggregation abilities of CAM, we have an oracle exploration where we explicitly say which rooms and items an agent has in its observations memory when answering the questions.
+Below we provide additional visualizations of the results of our learning-based CAM methods against MV and debating on other Matterport environments.
+Like Table 2, we use independent answers of $3$ LLMs explored with LGX.
+Majority of the CAM methods consistently outperform the baseline majority vote (MV) and debate methods, with Decision Tree (DT) and XGBoost (XG) being the highest-performing methods.
+Figure 4: 20 nodes
+Figure 5: 43 nodes
+Figure 6: 53 nodes
+Figure 7: 58 nodes
+Figure 8: 78 nodes
+Figure 9: 114 nodes
+Figure 10: 145 nodes
+Appendix E Decision Tree Visualization
+In this section, we showcase the potential interpretability of our learning-based CAM framework by visualizing the decision tree trained in Section Feature Importance Analysis.
 This tree in Figure 11 was trained on queries only pertaining to the kitchen and was trained on independent answers from four different LLM agents.
@@
 We provide the object encoding labels in Table 6.
-Figure 13: Visualization of Decision Tree from Section 7.
+Figure 11: Visualization of Decision Tree from Section Feature Importance Analysis.
 Object Label Encoding $o$

### v5

@@
-Central Answer Modeling for an Embodied Multi-LLM System
+Multi-LLM QA with Embodied Exploration
 Abstract
-Embodied Question Answering (EQA) is an important problem, which involves an agent exploring the environment to answer user queries.
-In the existing literature, EQA has exclusively been studied in single-agent scenarios, where exploration can be time-consuming and costly.
-In this work, we consider EQA in a multi-agent framework involving multiple large language models (LLM) based agents independently answering queries about a household environment.
-To generate one answer for each query, we use the individual responses to train a Central Answer Model (CAM) that aggregates responses for a robust answer.
-While prior Question Answering (QA) work has used a central module based on answers from multiple LLM-based experts, we specifically look at applying this framework to embodied LLM-based agents that must physically explore the environment first to become experts on their given environment to answer questions.
-Our work is the first to utilize a central answer model framework with embodied agents that must rely on exploring an unknown environment.
-We set up a variation of EQA where instead of the agents exploring the environment after the question is asked, the agents first explore the environment for a set amount of time and then answer a set of queries.
-Using CAM, we observe a $46\%$ higher EQA accuracy when compared against aggregation methods for ensemble LLM, such as voting schemes and debates.
-CAM does not require any form of agent communication, alleviating it from the associated costs.
-We ablate CAM with various nonlinear (neural network, random forest, decision tree, XGBoost) and linear (logistic regression classifier, SVM) algorithms.
-We experiment in various topological graph environments and examine the case where one of the agents is malicious and purposes contributes responses it believes to be wrong.
-Finally, we present a feature importance analysis for CAM via permutation feature importance (PFI).
-We provide and code and query dataset for further research.
-Introduction
-Embodied Question Answering (EQA) is an important task where an agent explores an environment, such as a household, to answer questions based on its observations.
-It has gained considerable interest in the past few years due to its applications for in-home robots and personalized assistants (Das et al.
-2018; Gordon et al.
-2018; Duan et al.
-2022a; Sima et al.
-2023; Luo et al.
-2023; Dorbala et al.
-2024a; Zhu et al.
-2023).
-For a seamless interface with humans, robotic agents must be equipped with natural language understanding (Xu et al.
-2021; Sun et al.
-2024) and possess common-sense reasoning of the world.
-One form of common sense is having a relational understanding of household items and their usual place (i.e., a mug is highly likely to be in the kitchen or living room).
-Foundation models, specifically Large Language Models (LLMs) such as GPT (Achiam et al.
-2023) and LLama (Touvron et al.
-2023), have shown remarkable prowess in both natural language understanding and common sense reasoning.
-Recently, prior art has used these foundation models in EQA for query generation (Dorbala et al.
-2024a) and answering (Majumdar et al.
-2024).
+Large language models (LLMs) have grown in popularity due to their natural language interface and pre trained knowledge, leading to rapidly increasing success in question-answering (QA) tasks.
+More recently, multi-agent systems with LLM-based agents (Multi-LLM) have been utilized increasingly more for QA.
+In these scenarios, the models may each answer the question and reach a consensus or each model is specialized to answer different domain questions.
+However, most prior work dealing with Multi-LLM QA has focused on scenarios where the models are asked in a zero-shot manner or are given information sources to extract the answer.
+For question answering of an unknown environment, embodied exploration of the environment is first needed to answer the question.
+This skill is necessary for personalizing embodied AI to environments such as households.
+There is a lack of insight into whether a Multi-LLM system can handle question-answering based on observations from embodied exploration.
+In this work, we address this gap by investigating the use of Multi-Embodied LLM Explorers (MELE) for QA in an unknown environment.
+Multiple LLM-based agents independently explore and then answer queries about a household environment.
+We analyze different aggregation methods to generate a single, final answer for each query: debating, majority voting, and training a central answer module (CAM).
+Using CAM, we observe a $46\%$ higher accuracy compared against the other non-learning-based aggregation methods.
+We provide code and the query dataset for further research.
+Multi-LLM QA with Embodied Exploration
+Bhrij Patel University of Maryland, College Park                        Vishnu Sashank Dorbala University of Maryland, College Park Dinesh Manocha University of Maryland, College Park                        Amrit Singh Bedi University of Central Florida
+1 Introduction
+Large Language Models (LLMs) (Achiam et al., 2023) become more embedded into society, with one of the most popular tasks being question-answering (QA) (Rawal et al., 2024; Kamalloo et al., 2023).
+Further research has investigated utilizing an ensemble of LLM models (Multi-LLM) for the ability to discuss (Li et al., 2023; Chen et al., 2023b), answer questions from large data sources and route questions to models with specialized knowledge (Mixture of Experts) (Li et al., 2024).
+However, in these prior works on ensemble LLMs for QA, the questions are asked in a zero-shot manner or information is provided to the models.
+For QA of a physical environment, an embodied agent must first explore an environment, such as a household, to answer questions based on its observations.
+It has gained considerable interest in the past few years due to its applications for in-home robots and personalized assistants Das et al.
+(2018); Gordon et al.
+(2018); Duan et al.
+(2022a); Sima et al.
+(2023); Luo et al.
+(2023); Dorbala et al.
+(2024a); Zhu et al.
+(2023).
+Possessing zero-shot common-sense reasoning of relational understandings of household items is crucial for LLMs to perform embodied exploration in unknown environments Dorbala et al.
+(2023, 2024b); Majumdar et al.
+(2024).
+Figure 1: Overview of our MELE framework.
+LLM-based agents that have explored the dynamic household are independently asked a binary “Yes/No” question and must answer based on their observations.
+Only one of the agents has the correct answer (“No”) while the other two agents are wrong.
+The aggregator function then takes in these responses to decide a final response.
 In existing works, only a single agent explores the environment with a limited field of view, inducing low coverage and high exploration costs.
-In contrast, for an LLM-based embodied agent to succeed in EQA, efficient exploration of the environment is needed as the questions will center around the state of the environment (Das et al.
-2018; Yu et al.
-2019; Zhu et al.
-2023; Majumdar et al.
-2024).
-Prior work has studied the zero-shot capabilities of LLM-based agents to explore an unknown household environment (Dorbala, Mullen Jr, and Manocha 2023; Dorbala et al.
-2024b; Majumdar et al.
-2024).
-The rise of Embodied Artificial Intelligence (EAI) (Duan et al.
-2022b) brings up the possibility of co-habited spaces with multiple agents and humans working in conjunction (Puig et al.
-2023).
-In this context, we consider a similar co-inhabited setup, and propose using a team comprising of multiple LLM-based agents or a Multi-LLM system for EQA.
-A Multi-LLM system allows for simultaneous exploration of the environment in the same amount of time.
+In contrast, for an LLM-based embodied agent to succeed in EQA, efficient exploration of the environment is needed as the questions will center around the state of the environment Das et al.
+(2018); Yu et al.
+(2019); Zhu et al.
+(2023); Majumdar et al.
+(2024).
+The rise of Embodied Artificial Intelligence (EAI) Duan et al.
+(2022b) brings up the possibility of co-habited spaces with multiple agents and humans working in conjunction Puig et al.
+(2023).
+With this trend, we believe investigating the abilities of multiple embodied LLM agents to answer questions on their environment becomes increasingly crucial for the personalization of AI.
+We consider a similar co-inhabited setup and propose using a team comprising multiple embodied LLM-based explorers.
+An ensemble system of explorers allows for simultaneous exploration of the environment in the same amount of time.
+However, there are also challenges when introducing a multi-agent setup.
 A key challenge in a multi-agent setup is with conflicting responses when agents are asked the same query.
-Since LLMs are blackboxes capable of generating free-form natural language, their responses to the same query may significantly vary (Gal and Grosz 2022; Wang et al.
-2024).
+Since LLMs are blackboxes capable of generating free-form natural language, their responses to the same query may significantly vary Gal and Grosz (2022); Wang et al.
+(2024b).
 In an embodied setting, such responses are further skewed by partial observations due to limited exploration.
-Prior work in ensemble LLM methods tackle this issue using weighted sum or voting schemes (Li et al.
-2024; Wei et al.
-2022; Wang et al.
-2022; Yang et al.
-2024) and consensus-reaching debates between the LLM agents (Chen et al.
-2023a; Du et al.
-2023; Xiong et al.
-2023; Guo et al.
-2024a; Chan et al.
-2023; Li, Patel, and Du 2023).
-However, such methods can be vulnerable to scenarios where incorrect answers from malicious or poor-performing agents can influence the overall decision (Amayuelas et al.
-2024).
-Figure 1: Overview of our Multi-LLM framework.
-LLM-based agents that have explored the dynamic household are independently asked a binary “Yes/No” question and must answer based on their observations.
-Only one of the agents has the correct answer (“Yes”) while the other two agents are wrong due to either not exploring the relevant portions of the house enough or being malicious.
-The central network, however from supervised training, has learned to weigh the decision of the first agent more when given context about the item, room, and timestep.
-An ideal EQA model would learn to identify incorrect agents and avoid being influenced by their responses.
-Furthermore, we would like to ideally avoid the communication costs incurred by debate-based approaches.
-Main Results: To address the issues presented above, we apply a Multi-LLM system to solve this embodied task.
+Prior work in ensemble LLM methods tackle this issue using weighted sum or voting schemes Li et al.
+(2024); Wei et al.
+(2022); Wang et al.
+(2022); Yang et al.
+(2024) and consensus-reaching debates between the LLM agents Chen et al.
+(2023a); Du et al.
+(2023); Xiong et al.
+(2023); Guo et al.
+(2024a); Chan et al.
+(2023); Li et al.
+(2023).
+However, such methods can be vulnerable to scenarios where incorrect answers from malicious or poor-performing agents can influence the overall decision Amayuelas et al.
+(2024).
+Main Results: To investigate the use of ensemble LLM for QA with embodied exploration, we apply a system of Multi-Embodied LLM Explorers (MELE) to solve this embodied task.
 We instruct the agents to answer based on their observations a set of binary embodied questions about the household.
-We then train a Central Answer Model (CAM) on the answers of independent agents with supervised learning.
-The network directly outputs the final answer to a given question and does not require any communication between agents.
-Furthermore, our model learns which agents to rely on while producing its final answer, which reduces the end user’s effort in having to individually figure out the best agent response.
+We then use various aggregation methods to output a final answer based on the independent responses from the agents: majority voting, debating, and training a supervised central answer model (CAM).
+Majority voting and debating are common aggregation methods in existing literature and we propose investigating training CAM for more robustness to adversarial or incorrect agents.
 Our novel contributions include:
-  • Multi-LLM QA Framework with Embodied Navigation: We train a Central Answer Model (CAM) for Question Answering in an Embodied Multi-Agent setting.
-We depart from prior EQA setups where the agents explore after being asked a question.
-Rather we have the agents explore before being asked a set of question on their environment.
-CAM is a classifier that aggregates multiple agent responses to predict an answer.
-We generate queries for various Matterport 3D (Chang et al.
-2017) environments.
-We train CAM on the labeled query dataset with multiple machine learning methods and show that our best model can achieve up to 46% higher accuracy than a majority vote and debate aggregation methods.
-While prior work has used a central answer module to output a final answer from a mixture of answers from multiple LLM experts (Puerto, Şahin, and Gurevych 2023; Si et al.
-2023; Jiang, Ren, and Lin 2023), we are the first to apply the framework to embodied agents that must first explore an environment.
-  • Integration with Exploration Systems: We evaluate our multi-agent framework on data gathered using an LLM-based exploration method (Dorbala, Mullen Jr, and Manocha 2023) on multiple agents in Matterport3D, demonstrating our system can work in conjunction with SOTA exploration methods in unknown environments.
-  • Examining Agent and Feature Reliance: We present a feature importance analysis of CAM reliance on each independent agent using the metric, permutation feature importance (PFI) (Breiman 2001; Fisher, Rudin, and Dominici 2018).
+  • Multi-LLM QA Framework with Embodied Navigation: We use multiple independent embodied LLM agents that explore the environment individually generate a response and use various aggregator functions to predict a final response.
+[^1] Link to anonymous repository: https://anonymous.4open.science/r/CAM-embodied-llm-54CE/ We depart from prior Embodied Question-Answering (EQA) Das et al.
+(2018) setups where the agents explore after being asked a question.
+Rather we have the agents explore before being asked a set of questions on their environment.
+We generate queries for various Matterport 3D Chang et al.
+(2017) environments.
+We train CAM on the labeled query dataset with multiple machine learning methods and show that the best CAM-based method can achieve up to 46% higher accuracy than a majority vote and debate aggregation methods.
+While prior work has used a central answer module to output a final answer from a mixture of answers from multiple LLM experts Puerto et al.
+(2023); Si et al.
+(2023); Jiang et al.
+(2023), we are the first to apply the framework to embodied agents that must first explore an environment.
+  • Integration with Exploration Systems: We evaluate our multi-agent framework on data gathered using an LLM-based exploration method Dorbala et al.
+(2023) on multiple agents in Matterport3D, demonstrating our system can work in conjunction with SOTA exploration methods in unknown environments.
+  • Examining Agent and Feature Reliance: We present a feature importance analysis of CAM reliance on each independent agent using the metric, permutation feature importance (PFI) Breiman (2001); Fisher et al.
+(2018).
 We also look at the agreement between the individual agent and the final answer of the central answer model and the baselines.
 Similarly to Si et al.
-(2023), some of our methods for the central answer model are interpretable and provide a Decision Tree visualization of one our models.
-Reference Exploration Central Module Multi-Agent Feature Importance
-(Majumdar et al. 2024)
-(Puerto, Şahin, and Gurevych 2023)
-Our Approach
-Table 1: Table of our contributions and novelties compared to recent prior work.
-This work is the first to tackle Embodied EQA with LLM-based Multi-agent system and to provide a feature importance analysis on the framework.
-Related Works
-Embodied Question Answering and Exploration
+(2023), some of our methods for the central answer model are interpretable and provide a Decision Tree visualization of one of our models.
+2 Related Works
+2.1 Embodied Question Answering and Exploration
 Embodied Question Answering (EQA) was first introduced by Das et al.
-(2018) and has become a popular task for robotics (Gordon et al.
-2018; Gordon, Fox, and Farhadi 2019; Duan et al.
-2022a; Sima et al.
-2023; Wijmans et al.
-2019; Chen et al.
-2019; Luo et al.
-2019), where traditionally an agent is given a query such as “was there a plant in the living room?’ and explores its environment, in this case a house, to answer the question.
+(2018) and has become a popular task for robotics (Gordon et al., 2018, 2019; Duan et al., 2022a; Sima et al., 2023; Wijmans et al., 2019; Chen et al., 2019; Luo et al., 2019), where traditionally an agent is given a query such as “was there a plant in the living room?’ and explores its environment, in this case a house, to answer the question.
 Zhu et al.
 (2023) had alternated between an exploration phase and then a question phase and repeated the process until the query was answered.
-Further work has extended EQA to scenarios where the final answer revolves around multiple objects (Yu et al.
-2019), and where a target object is not explicitly stated in the question (Tan et al.
-2023).
+Further work has extended EQA to scenarios where the final answer revolves around multiple objects Yu et al.
+(2019), and where a target object is not explicitly stated in the question Tan et al.
+(2023).
 Tan et al.
@@
 (2024) designed the first open-vocabulary EQA dataset to test and benchmark foundation models on this task.
-Recently, Singh, Pavlakos, and Stamoulis (2024) studied the zero-shot LLM performance on visual QA with indoor house scenes.
+Recently, Singh et al.
+(2024) studied the zero-shot LLM performance on visual QA with indoor house scenes.
 However, the LLM is given the scenes without any exploration of the environment.
-We evaluate our Multi-LLM approach with a SOTA LLM-based exploration method for observation gathering introduced by Dorbala, Mullen Jr, and Manocha (2023).
+We evaluate our Multi-LLM approach with a SOTA LLM-based exploration method for observation gathering introduced by Dorbala et al.
+(2023).
 Note: In our work, we deviate from the traditional EQA approach.
-We first have an exploration gathering phase, then answer a set of questions on the same environment.
-We do so for two reasons: 1) to provide a setup to better compare against Multi-LLM QA baselines: majority voting and debating; 2) we believe this setup is better suited for the scenario where the user asks a question about the items in their homes and quickly needs an answer.
-Multi-LLM Systems and Experts
-Research in Multi-LLM systems have grown in popularity with the rise of foundation models for areas such as task planning (Chen et al.
-2023b), application development (Wu et al.
-2023), and chatbots (Duan et al.
-2023).
+We first have an exploration gathering phase, then answer a set of questions in the same environment.
+We do so for two reasons: 1) to provide a setup that better resembles prior work in LLM-based MoE QA (Li et al., 2024) 2) we believe this setup is better suited for the scenario where the user asks a question about the items in their homes and quickly needs an answer.
+2.2 Systems of Multiple LLM-based Agents
+Research in LLM-based systems have grown in popularity with the rise of foundation models for areas such as task planning Chen et al.
+(2023c), application development Wu et al.
+(2023), and chatbots Duan et al.
+(2023).
 Extensive studies have focused on their abilities to cooperate.
-Prior work has used multi-LLM debate to reach a consensus (Chan et al.
-2023; Du et al.
-2023; Guo et al.
-2024a; Chen, Saha, and Bansal 2023).
+Prior work has used debate between LLM models to reach a consensus (Chan et al., 2023; Du et al., 2023; Guo et al., 2024a; Chen et al., 2023b).
 Chen et al.
 (2023a) analyzes the consensus-seeking debate between multiple LLM agents with different personalities, and Amayuelas et al.
-(2024) investigates the susceptibility of a multi-LLM discussion to an adversarial agent that purposefully tries to lead an incorrect answer choice.
+(2024) investigates the susceptibility of a LLM-based discussion to an adversarial agent that purposefully tries to lead an incorrect answer choice.
 For embodied AI, Guo et al.
 (2024b) have shown that multiple LLM-based agents can cooperate in teams by designating a leader for increased task efficiency.
-Other work have looked into using voting metrics of independent LLM-agents to arrive at a final answer (Li et al.
-2024; Wei et al.
-2022; Wang et al.
-2022).
+Other work have looked into using voting metrics of independent LLM-agents to arrive at a final answer (Li et al., 2024; Wei et al., 2022; Wang et al., 2022).
 Our work aims to remove the need for communication by training a central classifier on independent answers from multiple agents.
 Training a classifier in a supervised manner helps our approach become robust to unreliable agents that may influence overall answers in voting schemes.
-A central architecture to output a final answer from multiple answers from different LLM experts has been explored previously (Puerto, Şahin, and Gurevych 2023; Si et al.
-2023; Jiang, Ren, and Lin 2023).
+A central architecture to output a final answer from multiple answers from different LLM experts has been explored previously Puerto et al.
+(2023); Si et al.
+(2023); Jiang et al.
+(2023).
 Our approach differs from these prior works as the various LLM agents we utilize become experts in the environments from an embodied exploration phase and we show it can be used with off-the-shelf exploration methods.
-We believe that our approach can be considered a variation of Mixture-of-Experts (MoE) (Eigen, Ranzato, and Sutskever 2013), as early MoE methods relied on training ML models like SVM to learn the weightings of different experts, as our CAM framework does.
-Our main difference is that the MoE module also takes into account answers from embodied agents explore the environment.
-Figure 2: Overview of experimentation pipeline.
-[LEFT] We have an LLM answer a collection of embodied questions about the Matterport3D environment using the observations collected and stored by a specific agent during an exploration phase.
-The answers from the LLM are stored offline.
-[RIGHT] In the central network training phase, we select multiple sets of answers from different agents to train the classifier with the ground-truth answer with supervised learning.
-Multi-LLM System with Central Answer Model (CAM) for EQA
-Problem Formulation
+Furthermore, Wang et al.
+(2024a) introduced the idea of a Mixture of Agents to enhance task performance of LLMs.
+We specifically look at a mixture of embodied agents that explore a unknown environment.
+3 Multi-Embodied LLM Explorers
+3.1 Problem Formulation
 In this section, we provide a mathematical formulation of our specific EQA setting to characterize the EQA problem we are tackling in this work.
@@
 Let $s_{k}^{(o,r)}\in\{0,1\}$ be the independent answer to query $(o,r)$ of the $k$-th agent.
-Let $f$ be a function such that $f\left(o,r,\cup_{k=1}^{K}s_{k}^{(o,r)}\right)=\hat{y}\in\{0,1\}$.
-Given the query context $(o,r)$ and the independent agent answers to the context, $f$ should produce a final binary prediction to whether or not object $o$ is in room $r$.
-In this work, we present CAM as a viable method of the function $f$ over baselines majority voting and debating.
-We now describe CAM in the next section.
-Our Approach
-We implement a two-step procedure that comprises a question-answering phase, and a central model training phase.
+Let $f$ be a aggregation function such that $f\left(o,r,\cup_{k=1}^{K}s_{k}^{(o,r)}\right)=\hat{y}\in\{0,1\}$.
+Given the query context $(o,r)$ and the independent agent answers to the context, $f$ should produce a final binary prediction of whether object $o$ is in room $r$.
+3.2 Our Approach
+We implement MELE as a two-step procedure that comprises a question-answering phase, and a aggregation phase.
 The question-answering phase is an agent-specific step while the classification-training step relies on combining the intermediate results of the question-answering phases from various agents.
 Before starting the question-answering phase, the agents explored the environment to gather observations.
-Later, we examine the performance of our approach when using a LLM-based SOTA method for exploration.
-We describe both phases below.
-Question-Answering Phase
+Later, we examine the performance of MELE when using a LLM-based SOTA method for exploration.
+3.3 Question-Answering Phase
 In this phase, we generate a dataset of binary embodied questions specific to a Matterport 3D environment.
@@
 We use a topological graph environment because we are only concerned with high-level exploration policy in the house environment.
-Our approach can be used with other similar environments such as HM3D (Ramakrishnan et al.
-2021).
+Our approach can be used with other similar environments such as HM3D Ramakrishnan et al.
+(2021).
 Each LLM-based agent’s answers are based on independent observations from its exploration of the environment.
@@
 The observations are: [OBSERVATION DICTIONARY]
-Central Model Training Phase
-To give the final answer to each question, we train a model $f$ to give a final, binary “Yes/No” output.
-The model input dimension is $2+K$, where $K$ is the number of agents.
-The first two inputs account for the vectorized question in the form of “[object, room]”, $(o,r)$.
-The rest of the $K$ inputs come from entering the independent answers from each of the agents chosen for the multi-LLM system, $\cup_{k=1}^{K}s_{k}^{(o,r)}$.
-$f$ then outputs a predicted answer $\hat{y}$ and is then optimized with the ground truth answer $y$.
-During inference, we input unseen questions and agent answers into the network to output a test-time answer.
-Due to the modularity of our approach, for each set of independent responses, we train multiple machine learning models to produce multiple central answer models to compare against baseline aggregation approaches.
-Note: The classifier is overfitted to a given environment because the Multi-LLM embodied system is only for a specific environment.
-With our EQA variation, the agent must answer questions on what it has explored in the given environment.
-This overfitting allows for the personalization of our systems to the household environment.
-In our potential use case of a user asking a question of whether or not an object is located in a certain area of their environment, the agents should be trained and asked in the same environment.
-Experimentation
-We now describe our experimentation in detail.
-With these experiments, we aim to 1) measure the accuracy of CAM for EQA and show its performance over other aggregation baselines (Results section); 2) and 3) provide a feature importance analysis to highlight the benefit of utilizing independent agent responses in our approach (Feature Importance Analysis Section) In all experiments we use GPT-4-turbo.
-Query Generation
-We extract the embodied question bank from the Matterport3D environments.
-For each scan, we extract an offline list of observed items detected in each image of the environment using the GLIP object detection algorithm (Li et al.
-2022).
-The items space observed by GLIP can be found in the Appendix in Table 6.
-For each item and room combination, we formulate the question using the template described in Section Question-Answering Phase.
-Because these questions were extracted based on the observed items in the scan, we know the ground truth answer is “Yes”.
-To create questions with ground-truth answers “No”, for each query generated with a “Yes” ground-truth, we formulate a question with the same item but with a random room that is not within the subset of rooms that contains that specific item.
-Note: Our setup is highly dependent on the object detector used.
-Error can thus be introduced by the object detection with GLIP and then the LLM answer generation.
-Some outputs of GLIP included toilets in the living room and a TV in the hallway.
-When the LLM agent has a (room, item) in its observation memory in the system prompt, it will usually say Yes if asked about said room and item, but when it does not have it in memory, the questions are harder to answer, leading to error in answers generated by the independent LLM agent.
-More error analysis on agent answers will be studied in further in the Feature Importance Analysis section.
-Evaluation
-We use a SOTA algorithm, LGX (Dorbala, Mullen Jr, and Manocha 2023) for the exploration phase.
-Metric. To measure and compare the performances of our CAM approach against baseline aggregation approaches, we calculate the inference time accuracy, or correctness, of the model:
-$\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}_{y_{i}=\hat{y}_{i}},$ (1)
-where $y_{i}$ is the ground truth label of the test question, and $\hat{y}_{i}$ is the final answer prediction from CAM.
-We produce CAM models with various machine learning algorithms: Neural Network (NN), Random Forest (RF), Decision Tree (DT), XGBoost (XG), RBF SVM (SVM), Linear SVM (SVM-L), Logistic Regression (LR).
-The NN was implemented with Pytorch using 3 linear layers.
-The dimension were 16, 8, 1 in that order with a Sigmoid activation function at the end to ensure the output was in between 0 and 1.
-XGBoost used the default settings from the xgboost library package.
-RF from the sklearn library used $1000$ trees, and all other methods used the default settings from sklearn.
-Baselines
-We compare these different variations of CAM against two other forms of the aggregation methods used for ensemble LLM methods.
-Neither method is learning-based and requires no training set.
-We describe them below:
+3.4 Aggregation Phase
+We now describe the different methods used for aggregation function $f$.
+3.4.1 Non-learning-based Aggregation
+We first describe the non-learning based methods used commonly in ensemble LLM literature.
   • Majority Vote (MV): For each test question, we take the majority vote of the independent answers from the agents of the Multi-LLM system.
@@
   • Debate: For each test question, we have the agents of the Multi-LLM system participate in a turn-based discussion where they relay their thoughts to one another for 2 rounds.
-To initial the debate we provide each LLM-based agent context with the following system prompt.
+To initialize the debate we provide each LLM-based agent context with the following system prompt.
     System prompt: You are an embodied agent in a house.
@@
     User prompt: It is your turn to speak. Use your observation and conversation history to help.
-    When prompted for the final answer at the end of the debate. The user prompt is Please give your final ”Yes/No” answer. Give a definite answer.
+    When prompted for the final answer at the end of the debate.
+The user prompt is Please give your final "Yes/No" answer.
+Give a definite answer.
 We then collect the final answers from each agent and take the majority vote as the final prediction.
-Language-Guided Exploration
-Figure 3: LGX + CAM; Performance of CAM with $3$ independent agent that performed LGX for collecting observations.
-We use a 90%-10% train-test random split.
-We experiment on a Matterport environment with 215 nodes.
-We report the mean and standard deviation over $5$ trials.
-Against both baselines, all CAM methods outperform during test-time inference, with XGBoost having the highest accuracy.
+3.4.2 Central Answer Model (CAM)
+To give the final answer to each question, we train a model $f$ to give a final, binary “Yes/No" output.
+The model input dimension is $2+K$, where $K$ is the number of agents.
+The first two inputs account for the vectorized question in the form of “[object, room]”, $(o,r)$.
+The rest of the $K$ inputs come from entering the independent answers from each of the agents chosen for the multi-LLM system, $\cup_{k=1}^{K}s_{k}^{(o,r)}$.
+$f$ then outputs a predicted answer $\hat{y}$ and is then optimized with the ground truth answer $y$.
+During inference, we input unseen questions and agent answers into the network to output a test-time answer.
+Due to the modularity of our approach, for each set of independent responses, we train multiple machine learning models to produce multiple central answer models to compare against baseline aggregation approaches.
+Note: The classifier is overfitted to a given environment because the Multi-LLM embodied system is only for a specific environment.
+With our EQA variation, the agent must answer questions on what it has explored in the given environment.
+This overfitting allows for the personalization of our systems to the household environment.
+In our potential use case of a user asking a question of whether or not an object is located in a certain area of their environment, the agents should be trained and asked in the same environment.
+We produce CAM models with various machine-learning algorithms: Neural Network (NN), Random Forest (RF), Decision Tree (DT), XGBoost (XG), RBF SVM (SVM), Linear SVM (SVM-L), and Logistic Regression (LR).
+The NN was implemented with Pytorch using 3 linear layers.
+The dimension were 16, 8, 1 in that order with a Sigmoid activation function at the end to ensure the output was in between 0 and 1.
+XGBoost used the default settings from the xgboost library package.
+RF from the sklearn library used $1000$ trees, and all other methods used the default sklearn settings.
+4 Experimentation
+We now describe our experimentation in detail.
+With these experiments, we aim to 1) measure the accuracy of CAM for EQA and show its performance over other aggregation baselines (Results section); 2) and 3) provide a feature importance analysis to highlight the benefit of utilizing independent agent responses in our approach (Feature Importance Analysis Section).
+In all experiments we use GPT-4-turbo.
+4.1 Query Generation
+We extract the embodied question bank from the Matterport3D environments.
+For each scan, we extract an offline list of observed items detected in each image of the environment using the GLIP object detection algorithm Li et al.
+(2022).
+The items space observed by GLIP can be found in the Appendix in Table 5.
+For each item and room combination, we formulate the question using the template described in Section 3.3.
+Because these questions were extracted based on the observed items in the scan, we know the ground truth answer is “Yes".
+To create questions with ground-truth answers “No", for each query generated with a “Yes" ground-truth, we formulate a question with the same item but with a random room that is not within the subset of rooms that contains that specific item.
+Note: Our setup is highly dependent on the object detector used.
+Error can thus be introduced by the object detection with GLIP and then the LLM answer generation.
+Some outputs of GLIP included toilets in the living room and a TV in the hallway.
+When the LLM agent has a (room, item) in its observation memory in the system prompt, it will usually say Yes if asked about said room and item, but when it does not have it in memory, the questions are harder to answer, leading to error in answers generated by the independent LLM agent.
+More error analysis on agent answers will be studied further in the Section 5.
+4.2 Evaluation
+We use a SOTA algorithm, LGX Dorbala et al. (2023) for the exploration phase.
+Metric. To measure and compare the performances of our various MELE aggregation methods, we calculate the inference time accuracy, or correctness, of the model:
+$\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}_{y_{i}=\hat{y}_{i}},$ (1)
+where $y_{i}$ is the ground truth label of the test question, and $\hat{y}_{i}$ is the final answer prediction from the given aggregation method.
+4.3 Language-Guided Exploration
 Accuracy (%)
@@
 215 64.33 73.67 79.67 88.33 69.67 56.33 56.33 43.67 43.00
-Table 2: Performance measured in accuracy (Eq 1) of CAM with $3$ independent agent that performed LGX for collecting observations.
+Table 1: Performance measured in accuracy (Eq 1) of MELE with $3$ independent agent that performed LGX for collecting observations.
 We use a 90%-10% train-test random split.
@@
 In cases where one agent is malicious, our CAM method is still better suited than MV and debate.
+Figure 2: Performance of MELE methods with $3$ independent agent that performed LGX for collecting observations.
+We use a 90%-10% train-test random split.
+We experiment on a Matterport environment with 215 nodes.
+We report the mean and standard deviation over $5$ trials.
+Against both baselines, all CAM methods outperform during test-time inference, with XGBoost having the highest accuracy.
 We have all the LLM-based agents explore the household environment for a set amount of time before answering questions.
-In the exploration phase, each subordinate agent is set in the Matterport house scan graph and performs an augmented variation of Language-Guided Exploration (LGX) as proposed by Dorbala, Mullen Jr, and Manocha (2023).
+In the exploration phase, each subordinate agent is set in the Matterport house scan graph and performs an augmented variation of Language-Guided Exploration (LGX) as proposed by Dorbala et al.
+(2023).
 In traditional LGX, the LLM agent observes the immediate household items around itself and chooses which one to navigate to find a specific target item, forming a chain of “hops” to the item.
@@
 We have the $3$ agents perform $10$ steps of this augmented LGX and store the observations.
-The rest of the experiment process follows the same setup as described in the Approach Section with the Question-Answering and Central Model Training phases.
+The rest of the experiment process follows the same setup as described in Section 3 with the Question-Answering and Aggregation phases.
 We use a $90\%$-$10\%$ random train-test split over $5$ seeds (0 to 4 are used as the seeds).
-Experiments are ran on an Apple M1 Pro and macOS 14.5.
-Results: The performance of the various CAM methods and baselines in the $8$ graph environments is shown in Table 2.
+Experiments were run on an Apple M1 Pro and macOS 14.5.
+Results: The performance of the various CAM methods and non-learning aggregation methods in the $8$ graph environments is shown in Table 1.
 We test in environments of different sizes (measured in number of nodes) to showcase the generalization of the CAM approach across different size complexities.
-From the table, we see CAM model consistently outperforms both baselines in accuracy (Eq 1.
+From the table, we see CAM model consistently outperforms both MV and debating in accuracy (Eq 1.
 We report the average accuracy over $5$ seeds.
 We see that the non-linear methods consistently achieve higher accuracy than the linear methods, suggesting that the dataset contains no hyperplane separating the queries with “Yes” and “No” ground-truth answers.
-Specifically, XGBoost attains the highest accuracy in all but one environment, suggesting that the boosted ensemble method can best fit the generated query-answer data.
+Specifically, XGBoost attains the highest accuracy in all but one environment, suggesting that the boosted ensemble method best fits the generated query-answer data.
 Interestingly, in all but one of the environments, debating does slightly worse on average than MV.
@@
 This low agreement could stem from the errors introduced by the object detection of GLIP and the question-answering of each individual agent.
-We further analyze the independent answers impact on training in the Feature Importance Analysis Section.
+We further analyze the independent answers’ impact on training in the Section 5.
 Query Subject $\downarrow$ Object Kitchen Dining Room Bedroom Office
@@
 To turn the agent malicious we invert its independent answers so that all “Yes” answers turn to “No” and vice versa.
-We then train all CAM methods and execute baselines regularly.
+We then train all CAM methods and execute MV and debating regularly.
 Thus, the malicious agent contributes an answer it believes to be incorrect.
-From Table 2, we see that our CAM methods still outperform the MV and debate baselines.
+From Table 1, we see that our CAM methods still outperform MV and debate.
 We see that the debate baseline again does worse the majority vote, showing that the incorrect agent was able to sway the overall decision that was initially going to be correct.
-In the next section, we provide a feature importance of CAM.
-Feature Importance Analysis
+5 Feature Importance Analysis
 In the context of machine learning, feature importance is a metric by which input impacts the model’s final answer.
-For our Multi-LLM system, feature importance can be used to examine if any agent response or query context (i.e.
+For our MELE system, feature importance can be used to examine if any agent response or query context (i.e.
 room or object) is relied heavily upon for training the central answer model.
-To examine which features hold the most weight in the final decision of CAM model, we look into studying the variable, or feature importance (Breiman 2001; Fisher, Rudin, and Dominici 2018), of the CAM model.
+To examine which features hold the most weight in the final decision of the CAM model, we look into studying the variable, or feature importance Breiman (2001); Fisher et al.
+(2018), of the CAM model.
 We measure feature importance using permutation feature importance (PFI).
@@
 Please see the Appendix for a visualization of the Decision Tree to highlight the benefits of the interpretable ML-based CAM methods.
-Conclusion and Limitations
-In this paper, we tackle a variation of Embodied Question Answering with a Multi-LLM framework that aggregates the independent responses of LLM-based agents and produces a final answer.
-While other methods have used a central module to aggregate independent agent answers for QA tasks, we specifically investigate this method in the context of embodied exploration in a household environment.
-We propose a central answer model (CAM) that learns to weigh the responses of each agent to decide on a single answer to the given query.
-We show that various machine learning methods are suitable to train CAM and consistently outperform non-learning-based aggregation methods for ensemble outputs, specifically majority voting and multi-LLM debating.
-A key benefit of our system is that it requires no communication during inference, reducing time costs.
-Furthermore, our approach can be combined with SOTA methods for exploration in unknown households.
-We also show that CAM is more robust to malicious agents than the majority vote and debate baslines.
+6 Conclusion
+In this paper, we use Multi-Embodied LLM Explorers (MELE) to understand the effects of embodied exploration of multiple independent LLM-based agents on QA task for unknown environment.
+While other methods in the field of Mixture of Experts have used multiple LLM-based models for QA tasks, we specifically investigate this paradigm with a dependence on embodied exploration in a household environment.
+We compare the accuracy of various aggregation methods: majority voting, debating, and CAM.
+We show that various machine learning methods are suitable to train CAM and consistently outperform the non-learning-based aggregation methods.
+A key benefit of CAM is that it requires no communication during inference, reducing time costs.
+We also show that CAM is more robust to malicious agents than the majority vote and debate methods.
 We end this work by providing a feature importance analysis to study the effect of individual agents and target object specification on the final answer.
 We hope that this work will encourage research in embodied Multi-LLM for QA as LLM-based agents become more household.
+7 Limitations and Future Work
 This work has several limitations.
@@
 To better enhance practicality, future work should look into aggregation methods for situational, subjective, and non-binary questions.
-Appendix A Appendix
-Appendix B Code and Dataset
+It would also be useful to remove the need to explicitly mention the target object in the query.
+Other areas directions for research could be then to apply our CAM aggregation method to question-answering tasks outside of Embodied AI such as long video understanding.
+Third, our results hinge on the quality of the observation data.
+In our work, we relied on the GLIP model.
+Future work should analyze the effect of scene understanding algorithms on our setup.
+Appendix A Code and Dataset
 We provide an anonymous link to our code  [^2] https://anonymous.4open.science/r/CAM-embodied-llm-54CE/. The query dataset extracted is also within the submitted supplementary materials folder.
-Appendix C Multi-LLM vs LLM
+Appendix B Multi-LLM vs LLM
 To motivate multiple LLM agents we provide a comparison between 1 agent and 3 agents in Table 4.
@@
 Table 4: Rise of accuracy with multiple dynamic agents
-Appendix D Additional Visualization of Results
+Appendix C Additional Visualization of Results
 In this section, to focus on the aggregation abilities of CAM, we have an oracle exploration where we explicitly say which rooms and items an agent has in its observations memory when answering the questions.
@@
 Figure 5: 53 nodes
-Figure 7: 58 nodes
 Figure 6: 78 nodes
@@
 Figure 8: 145 nodes
-Appendix E Decision Tree Visualization
-In this section, we showcase the potential interpretability of our learning-based CAM framework by visualizing the decision tree trained in Section Feature Importance Analysis.
+Figure 9: Visualization of Decision Tree from Section 5.
+Appendix D Decision Tree Visualization
+In this section, we showcase the potential interpretability of our learning-based CAM framework by visualizing the decision tree trained in Section 5.
 This tree in Figure 9 was trained on queries only pertaining to the kitchen and was trained on independent answers from four different LLM agents.
@@
 We provide the object encoding labels in Table 5.
-Figure 11: Visualization of Decision Tree from Section Feature Importance Analysis.
 Object Label Encoding $o$